### v1

# Constraint Learning for Parametric Point Cloud

Abstract
Parametric point clouds are sampled from CAD shapes, have become increasingly prevalent in industrial manufacturing. However, most existing point cloud learning methods focus on the geometric features, such as local and global features or developing efficient convolution operations, overlooking the important attribute of constraints inherent in CAD shapes, which limits these methods’ ability to fully comprehend CAD shapes. To address this issue, we analyzed the effect of constraints, and proposed its deep learning-friendly representation, after that, the Constraint Feature Learning Network (CstNet) is developed to extract and leverage constraints. Our CstNet includes two stages. The Stage 1 extracts constraints from B-Rep data or point cloud. The Stage 2 leverages coordinates and constraints to enhance the comprehend of CAD shapes. Additionally, we built up the Parametric 20,000 Multi-modal Dataset for the scarcity of labeled B-Rep datasets. Experiments demonstrate that our CstNet achieved state-of-the-art performance on both public and proposed CAD shapes datasets. To the best of our knowledge, CstNet is the first constraint-based learning method tailored for CAD shapes analysis.
Project page: https://cstnetwork.github.io/
1 Introduction
Parametric point clouds are derived from parametric templates, which consist of primitives, constraints, and dimensional parameters. By specifying all parameters, the templates can be instantiated into CAD shapes, as illustrated in Figure 1. Parametric point clouds provide detailed information that is essential for various engineering applications, such as model design and product machining.
In recent years, deep learning methods have emerged as powerful tools for point cloud analysis, with models such as PointNet [32], PointNet++ [33], and PointCNN [24] demonstrating significant success across various tasks. These models are primarily designed and evaluated on graphic objects, such as animals, trees, and other free-form shapes, where shape comprehension relies on geometric features visible in contours. However, some CAD shapes, despite exhibiting minimal visual differences, serve distinct purposes due to variations in their functional regions. This similarity in appearance poses challenge in distinguishing them based on geometric features. Notably, variations in the functional regions often introduce differing constraints, as these regions need to fit closely with other parts to perform functions such as transmission and positioning. Thus, distinguishing them from constraints-based perspective becomes more feasible, as illustrated in Figure 1.
Figure 1: Motivation. Some similar CAD shapes serve different functions. The eccentric wheel has a rotation axis that is offset from its outer cylindrical surface, facilitating precise positioning or tension adjustment. In contrast, the flywheel has a coaxial rotation axis with its outer cylindrical surface and is used to store rotational kinetic energy. Although these shapes are challenging to distinguish visually, they can be easily differentiated from the perspective of constraints.
Additionally, the limited availability of B-Rep datasets poses barrier to deep learning on CAD shapes. Most existing CAD shapes datasets consist of mesh files, such as MCB [16] and ESB [12]. While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information. In contrast, B-Rep data serves as the native representation of CAD shapes and is therefore more suitable for datasets construction. However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files. Although the ABC [18] contains a large number of B-Rep files, it remains unlabeled.
To facilitate the constraint feature learning of CAD shapes, we conducted studies on deep learning methods and dataset development. Traditional constraint definitions are not well-suited for deep learning, as they vary across constraint types and involve multiple possible combinations to represent primitives’ relation. To address this challenge, we proposed a novel constraint representation as point-wise local attributes. Afterwards, we developed the CstNet for constraint extraction and feature learning, which consist of two stages. The stage 1 is built for constraint acquisition, if B-Rep data is avilable, the openCASCADE Technology (OCCT) is used to extracts constraints, otherwise the constraint prediction from point cloud (CST-PCD) module is adopted. The CST-PCD is designed relies on local features only, enabling it to generalize to unseen datasets after pre-training. The stage 2 leverages both point cloud and constraints for constraint feature learning, which employs both point-level and feature-level attention for effective feature extraction, enhancing the comprehension of CAD shapes. Finally, we built up a multi-modal classification dataset Parametric 20000, which contains B-Rep, Mesh, and point cloud. Given the current scarcity of labeled CAD shapes datasets, Parametric 20000 offers researchers an expanded resource for their work. Our contributions are threefold:
  • We introduce the deep learning friendly expression of constraints.
  • We designed the CstNet for constraint extraction and parametric feature learning.
  • We collected and curated the Parametric 20000 dataset, which consist of mesh, point cloud, and B-Rep data.
Our CstNet has been validated across various tasks, on the mechanical classification dataset MCB [16], our model achieved an overall accuracy improvement of 2.98 % compared to the state-of-the-art, and for mechanical segmentation dataset 360 Gallery [19], that is 6.74 %.
2 Related Work
For the published deep learning methods focus on CAD shape are very few, and even fewer aimed at parametric point cloud, this section will provide an overview of existing point cloud learning methods as well as CAD shapes oriented models. By integrating insights from above areas, a comprehensive review of the relevant works could be presented.
2.1 Deep learning on point clouds
Point cloud deep learning can be categorized into three types based on data processing approaches: view-based, voxel-based, and point-based methods.
View-based methods typically project the point cloud from multiple views, the obtained figures are then fed in convolutional neural networks (CNNs) for feature extraction [37, 15, 6, 54, 53, 42]. However, these methods face significant challenges when applied to large-scale scanned data, primarily due to the difficulty of covering entire scenes from single-point perspectives, and the choice of projection views significantly impacts recognition performance. Besides that, it cannot be extended to 3D data segmentation or reconstruction easily.
Voxelization-based methods convert point cloud into 3D voxels, and then processed by 3D CNNs for feature extraction [30, 47, 35, 40, 17, 20]. However, the size and orientation of the voxels can affect computational efficiency, memory usage. The voxel grid quantization process may lead to geometric detail lost, and the inherent sparsity of point clouds may result in many useless voxels.
Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [32, 33, 24, 10, 52, 2, 38, 27, 55, 56, 9, 7, 50, 49, 48, 34, 28, 26, 41, 46, 4]. Typically, symmetric functions and k-nearest neighbors (KNN) are employed for feature extraction. However, the unordered nature and transformation invariance of point clouds pose challenges for efficient feature extraction.
2.2 Deep learning on CAD shapes
CAD shapes deep learning can be categorized into three types based on the input data format: B-Rep-based learning, mesh-based learning, and sketch-based learning.
B-Rep-based methods typically involve B-Rep data conversion and feature extraction. Since B-Rep data cannot be directly fed into deep learning models, the challenge lies in converting B-Rep into deep learning compatible format, with graph representation is commonly used. These approaches are commonly applied to classification [13], segmentation [19, 5], assembly [14, 43], engineering sketch generation [11], and CAD operation sequence prediction [45, 57].
Mesh-based learning methods use triangle models as input, such as STL and OBJ. Compared to B-Rep, mesh representations lack boundary details, which prevents these methods from leveraging data conversion for improved performance. As a result, contributions often rely on mechanical datasets provided by the authors. These methods are primarily used for classification [29], retrieval [3], segmentation [39], and B-Rep generation [8].
Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [8, 44, 31, 36, 25], 3D models [21, 22], or CAD operation sequences [51]. Most of these models are generative, and utilizing transformer modules frequently. A key challenge is incorporating the regularity and connectivity of primitives, which typically overlooked in graphic objects analysis.
While there has been substantial progress in above domains, the engineering applications of parametric point cloud analysis demand further innovation to ensure accuracy and robustness.
3 Method
3.1 Problem Statement
Constraints is crucial for CAD shapes, can they be utilized to enhance deep learning methods’ performance? To answer above question, we designed the following experiments. The objective is to compare the prisms shown in Figure 2 with cuboids. In each experiment, only prisms with a specific angle are selected, creating a binary classification task between prisms and cuboids. Consequently, eight independent experiments are conducted to cover all angles. As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them. The constraint-aware model is built on PointNet++ backbone, which predict constraints and then adopt it for classification, more details in supplementary.
Figure 2: Prisms compared to cuboid.
The experimental results are presented in Figure 3. For prisms with angle of 50°50°50  ° and 89°89°89  °, the constraints have little improvement on the results. The reason lies in the 50°50°50  ° prism differs substantially from the cuboid, while the 89°89°89  ° prism is too similar to the cuboid. For other circumstances, incorporating constraints improves classification accuracy significantly, demonstrating that the constraint enhanced model’s comprehension on CAD shapes.
Figure 3: Classification accuracy of validation experiments.
However, the constraint learning method designed in the validation experiments is only applicable to simple scenarios. To develop a more generalized approach, we further investigated the constraint representation and deep learning architecture tailored for parametric point clouds.
3.2 Constraint Representation
Constraint should be defined as point-wise attribute for point cloud, analogous to normal vector, as point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation.
Following the constraint definition in CAD software is not applicable, However, once constraint is applied, the positional relations between primitives become fixed. Therefore, if the type of primitives and their relations are determined, the constraint on CAD shape can be defined comprehensively.
The point-wise primitive type could be represented as the type of the primitive which the point aligned to, using one-hot encoding.
The positional relation is defined as the combination of main axis direction and adjacency. Considering that the direction of primitives’ main axis is unique, as shown in Figure 4, relations such as parallel and vertical could be derived by the direction of their main axes. However, some dimensional relations can not, such as the cylinders’ radius equivalence. Thus, additional data is required to supplement this information.
Figure 4: Main axis of primitives. Plane: normal vector. Cylinder, cone, and sphere: rotation axis.
Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitation on the distance and size of other primitives. For example, in a cuboid, the adjacency and normal of the six faces ensures that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one. Therefore, if the adjacency combined with main axis directions, the positional relation can be determined comprehensively.
Adjacency represents the connection between primitives and cannot be expressed as point-wise attribute directly. The edges are the key indicator of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges. For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by two-channel one-hot vector.
In summary, each composition of constraints in parametric point cloud is illustrated in Figure 5, and each point with constraint is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency).
Figure 5: Constraint representation of parametric point cloud.
3.3 Method Overview
Our CstNet and module details are depicted in the top and bottom side of Figure 6 respectively. It should be noted that the CST-PCD module requires pre-training, and once the pre-training is completed, its parameters are kept frozen. Given the locality of the constraints expression we designed, the pre-trained CST-PCD module is able to generalize to other unseen mechanical datasets directly. In this work, the dataset for pre-training CST-PCD module comes from ABC [18].
Figure 6: Overview of CstNet. The top side illustrates the overall architecture of CstNet, comprising two stages. Stage 1 is designed for constraint acquisition. When B-Rep data is available, the CST-BRep module is utilized to extract constraints; otherwise, use the CST-PCD predicts constraints. Stage 2 perform constraint feature learning, facilitate deeper understanding for CAD shapes. The bottom side presents the details of module design.
3.4 Constraint Acquisition
When B-Rep data is available, the CST-BRep module can be used to compute constraints.
For the input B-Rep data, tessellation is performed by OCCT to generate a mesh model, followed by Poisson disk sampling [23] via pymeshlab to ensure a uniform point distribution and control the point count, if point cloud generation is not needed, this step could be skipped.
For adjacency computation, we identify valid edges firstly. In this study, an edge is considered invalid if the two adjacent faces are smoothly connected along the edge, this work is accomplished by OCCT. Next, we compute the distance from each point to all valid edges. A point is marked as near an edge if the minimum distance is below a threshold (determined by the model surface area, set to $0.08\times Area/4\pi$ in this paper). For the primitive type and main axis direction of a point, we first determine which face it aligned with, OCCT is then used to analyze the face, providing both the face type and main axis direction.
In the absence of B-Rep data, such as working with mesh or voxel models, it can be converted into point cloud, after which the CST-PCD module can be used to predict constraints. Since each point’s constraints are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary. The constraint computation of a point $\boldsymbol{p}$ could be divided into the following steps:
Step 1. Identify neighbor points around $\boldsymbol{p}$.
Step 2. From all neighbor points, identify which belong to the same primitive as $\boldsymbol{p}$, these points are referred to as valid points.
Step 3. Fit shapes such as cylinders or planes using the valid points, and determine the primitive type with the smallest fitting error.
Step 4. Based on the the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
While the above process can be carried out using traditional methods, it is highly complex, which is why a deep learning model is employed to achieve the same.
The design of CST-PCD module is inspired by the aforementioned calculation process. For Step 1, which involves finding neighbor points, the traditional approach is to use the K Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere. However, for constraint computation, it is more effective to search for neighbors along the shape’s surface. This approach captures more relevant points while reducing the influence of redundant points, as shown in Figure 7. To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary.
Figure 7: Comparison of KNN and SurfaceKNN. Blue point: center, red point: neighbor.
For Step 2, which involves identifying valid points from neighbors, the attention mechanism between points is employed to approximate this process, which applying different weight to neighbor points when updating the center point’s embedding. We designed the point attention layer based on vector self-attention, the relation between points is designed as embedding subtraction, as shown in Equation (1).
$\boldsymbol{f}^{\prime}_{i}=\sum_{f_{j}\in\mathcal{N}_{i}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{i})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{i}).$ (1)
Where $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication. $\mathrm{Q}(\boldsymbol{f})$, $\mathrm{K}(\boldsymbol{f})$, and $\mathrm{V}(\boldsymbol{f})$ are defined in the same way by concatenating $\boldsymbol{f}$ with the positional encoding $\Delta$ and then feeding it into MLP, where $\Delta$ is defined as $\Delta=\boldsymbol{p}_{i}-\boldsymbol{p}_{j}$.
Step 3 and Step 4 involve determining the Primitive Type and Main Axis Direction, both steps can be accomplished by MLPs. We employed a U-Net-like structure to enhance information utilization. In the CST-PCD module, C-MLP refers to concatenating the input tensors and feeding them into MLPs. The final MLP in this module consists of three parallel MLPs with the same input, where the outputs correspond to Primitive Type, Main Axis Direction, and Adjacency.
The structure of CST-PCD module reveals that only local information is utilized, which aligns with the inherently local nature of the constraint representation we designed. For no global information is incorporated and most CAD shapes consist of a limited set of local geometry, such as plane, cylinder, and sphere, the CST-PCD can generalize to unseen datasets if pre-trained on a large dataset. As a result, our CST-PCD enables the effective use of existing unlabeled B-Rep datasets, making it possible to serve for meaningful tasks.
3.5 Constraint Feature Learning
The stage 2 of CstNet is designed to leverage constraints to deepen the model’s understanding of CAD shapes. The designed architecture extract features from Primitive Type, Main Axis Direction, and Adjacency separately, attention between features is then applied to update constraint feature. The structure and module details are illustrated in Figure 6.
The Triple MLP takes the point cloud, Primitive Type, Main Axis Direction, and Adjacency as inputs. In this module, the point cloud is concatenated with the Primitive Type, Main Axis Direction, and Adjacency separately before being fed into three individual MLPs. The outputs are the Primitive Feature, Axis Feature, and Adjacency Feature, respectively.
In the C-MLP layer, point cloud is concatenated with Primitive Type, Main Axis Direction, Adjacency, and then feed into MLP, with the output being the Parametric Feature.
[…]
The Fea Attention layer also takes the Primitive Feature, Axis Feature, Adjacency Feature, and Parametric Feature as input, attention between features is employed to update constraint feature efficiently. This design is considered that different points may focus on different features, for example, points near edge may focus more on the Adjacency Feature, while others may pay more attention to the Primitive or Axis Feature. The Fea Attention is developed by vector attention as formulated in Equation (2).
$\boldsymbol{f}^{\prime}_{cst}=\sum_{\boldsymbol{f}_{j}\in\mathcal{F}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{cst})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{cst}).$ (2)
In which =$\mathcal{F}=\{\boldsymbol{f}_{pmt},\boldsymbol{f}_{axs},\boldsymbol{f}_{adj},\boldsymbol{f}_{cst}\}$, $\boldsymbol{f}_{pmt}$ is Primitive Feature, $\boldsymbol{f}_{axs}$ is Axis Feature, $\boldsymbol{f}_{adj}$ is Adjacency Feature, and $\boldsymbol{f}_{cst}$ is Constraint Feature. $\Delta$ is defined as one-hot encoding of $\boldsymbol{f}_{j}$ type.
For downstream tasks, if classification is required, the final constraint feature can be fed into MLP and softmax. For segmentation tasks, upsampling and skip link concatenation similar to those in PointNet++ [33] can be applied.
4 Experiments
4.1 Experimental Settings
In general, we train CstNet using the Negative Log Likelihood Loss (NLLLoss) and optimized by Adam optimization for 200 epochs, with a weight decay of 0.0001 and an initial learning rate of 0.0001. We use StepLR learning rate scheduler, with step size of 20 and gamma of 0.7. The batch size is 16. All experiments are performed on GeForce RTX 4090 GPU.
Parametric 20000 dataset: our Parametric 20000 dataset contains 19,739 multi-modal instances categorized into 75 classes. Each instance includes B-Rep data, mesh, and point cloud. All B-Rep data are collected from the following three sources. More details in the supplementary.
Sources 1 (36.05 %): Downloaded and purified from TraceParts [16]. The CAD shapes on TraceParts are collected from manufacturing companies.
Sources 2 (24.12 %): Provided by a CAD company [^1] https://www.newdimchina.com/ that collaborates with our team. These shapes are all mechanical components.
Sources 3 (39.83 %): Instantiated from parametric templates designed by our team. These templates are designed based on common mechanical structures.
Our Parametric 20000 dataset encompasses various common formats of CAD shapes, making it suitable for a wide range of tasks. The careful selection and purification process undertaken by our team further ensures its high quality.
4.2 Constraint Acquisition
The extraction of constraints from B-Rep data is accomplished by CST-BRep module, with results visualized in the "GT" column of Figure 8. By referring the model row, this module performs well in extracting three components of constraint.
Extract constraints from point cloud is achieved by CST-PCD module. The training set is sourced from 25 Trunks of ABC [18], with each trunk containing 10,000 B-Rep files. We utilized CST-BRep to extract constraints as Ground Truth for model training. After the training finished, the prediction accuracy of Primitive Type reached 94.23 %, for Adjacency, that is 94.62 %, and the MSE loss for Main Axis Direction is 0.0559.
In subsequent classification and segmentation tasks, above pre-trained CST-PCD module is utilized to predict constraints with its weights frozen and no further updates. To ensure that only point cloud is used as input across all methods in comparison, even when B-Rep data is available, the CST-BRep module will not be activated to obtain constraints.
After CST-PCD module is pre-trained, we evaluated it on a new Trunk of ABC. The visualization of constraint prediction is shown in the "Pred." column of Figure 8. After traversing this Trunk, the prediction accuracy of Primitive Type reached 91.19 %, for Adjacency, that is 90.24 %, the MSE loss of Main Axis Direction is 0.0998, above demonstrating the strong capability for constraint prediction.
Figure 8: Constraint prediction on ABC. MAD: Main Axis Direction. PT: Primitive Type. ADJ: Adjacency (whether point is near edge). Main Axis Direction has been processed for visualization, details in supplementary.
Additionally, we evaluated the pre-trained CST-PCD module on the CAD shapes dataset MCB [16] to predict constraints. For the MCB consists of mesh files, it was not possible to extract ground truth constraints for comparison. Based on the results visualized in Figure 9, CST-PCD demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
Figure 9: Constraint prediction on MCB.
4.3 Classification
The classification results on the Parametric 20000 dataset are presented in Table 1, where our CstNet outperformed other methods across both metrics. Additionally, the instance accuracy of all methods is below 90 %, with a notable performance gap between them, indicating that this dataset poses challenges and effectively differentiates between methods.
Table 1: Classification results on Parametric 20000. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
Method Acc Acc* F1 mAP
PointCNN [24] 61.92 55.29 53.97 40.99
PointNet [32] 75.61 74.08 74.23 82.15
PointNet++ [33] 82.89 86.80 82.31 88.05
PointConv [46] 80.53 82.10 81.86 86.54
DGCNN [41] 85.99 88.20 87.20 89.20
3DGCN [26] 86.19 86.31 85.37 87.64
Ours 89.93 91.04 90.32 92.32
We conducted evaluations on the CAD shapes dataset MCB [16], with results presented in Table 2. Our CstNet achieved the highest scores across all metrics on MCB-A, but performed less effectively on MCB-B. This discrepancy may stem from the composition of MCB: MCB-A includes shapes from TraceParts (69.27 %), 3D Warehouse (21.70 %), and GrabCAD (9.03 %), while MCB-B is limited to shapes from 3D Warehouse and GrabCAD. TraceParts provides data from manufacturing companies, whereas 3D Warehouse and GrabCAD contain shapes for general usages [16]. As the CST-PCD module in CstNet is pre-trained on ABC, which primarily comprises mechanical components, this pre-training may lead to poor constraint prediction performance on MCB-B, thereby influencing the classification results.
Table 2: Classification results on MCB. A / B column: results of MCB A / B subset.
Method Acc.instance % Acc.class % F1-score mAP %
A B A B A B A B
PointCNN [24] 93.89 93.67 81.85 86.80 83.86 88.63 90.13 93.86
PointNet [32] 86.78 80.36 67.70 65.30 86.55 79.89 74.08 71.66
PointNet++ [33] 87.45 93.91 73.68 87.97 88.32 73.45 91.33 91.37
SpiderCNN [52] 93.59 89.31 79.70 79.29 81.30 80.72 86.64 82.47
PointConv [46] 93.25 90.64 80.24 80.17 71.31 76.07 82.19 86.09
DGCNN [41] 92.54 90.75 74.47 78.17 76.12 79.69 74.27 79.73
3DGCN [26] 93.71 92.13 78.71 84.31 84.59 79.24 84.35 84.35
Ours 96.87 92.61 89.21 85.94 89.85 85.68 93.17 91.25
4.4 Segmentation
We conducted experiments on the mechanical segmentation dataset 360 Gallery [19], which contains point clouds with segmentation labels based on CAD modeling operations. However, this labeling approach presents disadvantage for point cloud segmentation. For example, a cube in this dataset has four faces labeled as "ExtrudeSide" and two faces labeled as "ExtrudeEnd", which poses an unrealistic task for point cloud networks. Additionally, the distribution of points across different labels is highly imbalanced. For instance, points labeled as "ExtrudeSide" accounts for 45.94 %, while those labeled as "RevolveEnd" make up only 0.12 %. Despite these limitations, to our knowledge, 360 Gallery is currently the only publicly available point cloud dataset for CAD shapes segmentation task.
The segmentation results are presented in Table 3, where our method outperforms the state-of-the-art methods on most metrics. The results also highlight the limitations of this 360 Gallery in terms of point cloud segmentation. Apart from "ExtrudeSide" and "ExtrudeEnd", none of the models were able to differentiate the other segmentation categories except DGCNN managed to predict "RevolveSide" with a low mIoU.
Table 3: Part Segmentation Results on 360 Gallery. PN: PointNet. PN2: PointNet++. DGN: DGCNN. 3DN: 3DGCN. OA: overall accuracy %. mIoU: instance mIoU. mIoU*: class mIoU. ESide: extrude side mIoU, EEnd: extrude end mIoU. Other segmentation categories’ mIoU are all 0 except RevolveSide mIoU of DGCNN is 9.60, more datails in supplementary.
Method OA mIoU mIoU* ESide EEnd
PN [32] 42.13 18.89 7.04 19.83 36.47
PN2 [33] 58.18 26.98 11.27 49.80 40.36
DGN [41] 64.02 35.54 14.36 54.48 50.82
3DN [26] 60.58 33.03 12.23 50.67 47.20
Ours 70.76 46.65 20.51 61.29 60.65
A portion of the segmentation results on 360 Gallery test set are shown in Table 4, where our CstNet shows segmentation results closest to the Ground Truth. PointNet’s segmentation differs from other methods, for instance, in the third and fourth rows, PointNet mistakenly segments a single primitive into two different categories. This discrepancy likely arises from the model’s inability to effectively capture local features.
Table 4: Visualization of part segmentation on 360 Gallery. GT: ground truth. PN: PointNet [32]. PN2: PointNet++ [33]. DGN: DGCNN [41]. 3DN: 3DGCN [26].
GT PN PN2 DGN 3DN Ours
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
4.5 Ablation Studies
Our model leverages all three components of constraint: Primitive Type, Main Axis Direction, and Adjacency. Additionally, we utilized the SurfaceKNN to obtain more suitable neighbors for each point. To validate the effectiveness of our design, we conducted a series of experiments on Parametric 20000 dataset.
For each component of constraint, we evaluated their importance by disable one at a time. This operation is conducted by setting it as a zero tensor before feeding into the model. The results are presented in Table 5, which show that all three components contribute to improve the performance of our CstNet.
To assess the effectiveness of SurfaceKNN, we compared it with KNN. From Table 5, it can be concluded that in most cases, our SurfaceKNN outperforms KNN. However, when Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
Table 5: Effects on Parametric 20000 dateset with varying neighbor search and constraints. Row 3 - 6: SurfaceKNN, Row 7 - 10: KNN.
Constraint Acc Acc* F1 mAP
MAD ADJ PT
89.93 91.04 90.32 92.32
86.32 89.05 88.38 90.59
88.14 89.61 88.64 91.35
88.68 88.94 89.01 91.22
89.03 90.00 89.22 91.43
85.04 85.93 85.74 88.08
87.27 88.23 87.98 90.00
88.73 88.97 88.91 91.21
5 Conclusions
In this paper, we studied the CAD shapes and introduced the constraints expression which suitable for deep learning methods. Afterward, we proposed the CstNet for extracting and learning constraint features. Additionally, we built up a multimodal CAD shape classification dataset Parametric 20000. Experimental results on both public datasets and proposed dataset demonstrate that our CstNet achieves state-of-the-art performance across multiple metrics. Comprehensive ablation studies further validate that each component of constraint and SurfaceKNN contributes to improve the performance of CstNet.
CAD shapes possess numerous attributes. In this paper, we analyze the constraints from functional perspective. In future work, we plan to investigate additional aspects such as part design and manufacturing processes, to achieve a deeper understanding of CAD shapes.

### v2

@@
 Abstract
-Parametric point clouds are sampled from CAD shapes, have become increasingly prevalent in industrial manufacturing.
-However, most existing point cloud learning methods focus on the geometric features, such as local and global features or developing efficient convolution operations, overlooking the important attribute of constraints inherent in CAD shapes, which limits these methods’ ability to fully comprehend CAD shapes.
-To address this issue, we analyzed the effect of constraints, and proposed its deep learning-friendly representation, after that, the Constraint Feature Learning Network (CstNet) is developed to extract and leverage constraints.
+Parametric point clouds are sampled from CAD shapes, and have become increasingly prevalent in industrial manufacturing.
+However, most existing point cloud learning methods focus on the geometric features, such as developing efficient convolution operations, overlooking the important attribute of constraints inherent in CAD shapes, which limits these methods’ ability to comprehend CAD shapes fully.
+To address this issue, we analyzed the effect of constraints, and proposed its deep learning-friendly representation, after that, the Constraint Feature Learning Network (CstNet) was developed to extract and leverage constraints.
 Our CstNet includes two stages.
-The Stage 1 extracts constraints from B-Rep data or point cloud.
-The Stage 2 leverages coordinates and constraints to enhance the comprehend of CAD shapes.
+Stage 1 extracts constraints from B-Rep data or point cloud.
+Stage 2 leverages coordinates and constraints to enhance the comprehension of CAD shapes.
 Additionally, we built up the Parametric 20,000 Multi-modal Dataset for the scarcity of labeled B-Rep datasets.
-Experiments demonstrate that our CstNet achieved state-of-the-art performance on both public and proposed CAD shapes datasets.
-To the best of our knowledge, CstNet is the first constraint-based learning method tailored for CAD shapes analysis.
+Experiments demonstrate that our CstNet achieved state-of-the-art performance on both public and proposed CAD shape datasets.
+To the best of our knowledge, CstNet is the first constraint-based learning method tailored for CAD shape analysis.
 Project page: https://cstnetwork.github.io/
@@
 Parametric point clouds are derived from parametric templates, which consist of primitives, constraints, and dimensional parameters.
-By specifying all parameters, the templates can be instantiated into CAD shapes, as illustrated in Figure 1.
+By specifying all parameters, the templates can be instantiated into CAD shapes, as illustrated in Fig.
+1.
 Parametric point clouds provide detailed information that is essential for various engineering applications, such as model design and product machining.
@@
 However, some CAD shapes, despite exhibiting minimal visual differences, serve distinct purposes due to variations in their functional regions.
-This similarity in appearance poses challenge in distinguishing them based on geometric features.
+This similarity in appearance poses challenges in distinguishing them based on geometric features.
 Notably, variations in the functional regions often introduce differing constraints, as these regions need to fit closely with other parts to perform functions such as transmission and positioning.
-Thus, distinguishing them from constraints-based perspective becomes more feasible, as illustrated in Figure 1.
+Thus, distinguishing them from a constraint-based perspective becomes more feasible, as illustrated in Fig.
+1.
 Figure 1: Motivation.
@@
 Although these shapes are challenging to distinguish visually, they can be easily differentiated from the perspective of constraints.
-Additionally, the limited availability of B-Rep datasets poses barrier to deep learning on CAD shapes.
-Most existing CAD shapes datasets consist of mesh files, such as MCB [16] and ESB [12].
+Additionally, the limited availability of B-Rep datasets poses barriers to deep learning on CAD shapes.
+Most existing CAD shape datasets consist of mesh files, such as MCB [16] and ESB [12].
 While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information.
-In contrast, B-Rep data serves as the native representation of CAD shapes and is therefore more suitable for datasets construction.
+In contrast, B-Rep data [19, 56] serves as the native representation of CAD shapes and is therefore more suitable for dataset construction.
 However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files.
@@
 To facilitate the constraint feature learning of CAD shapes, we conducted studies on deep learning methods and dataset development.
-Traditional constraint definitions are not well-suited for deep learning, as they vary across constraint types and involve multiple possible combinations to represent primitives’ relation.
+Traditional constraint definitions are not well-suited for deep learning, as they vary across constraint types and involve multiple possible combinations to represent primitives’ relations.
 To address this challenge, we proposed a novel constraint representation as point-wise local attributes.
-Afterwards, we developed the CstNet for constraint extraction and feature learning, which consist of two stages.
-The stage 1 is built for constraint acquisition, if B-Rep data is avilable, the openCASCADE Technology (OCCT) is used to extracts constraints, otherwise the constraint prediction from point cloud (CST-PCD) module is adopted.
-The CST-PCD is designed relies on local features only, enabling it to generalize to unseen datasets after pre-training.
-The stage 2 leverages both point cloud and constraints for constraint feature learning, which employs both point-level and feature-level attention for effective feature extraction, enhancing the comprehension of CAD shapes.
+Afterward, we developed the CstNet for constraint extraction and feature learning, which consist of two stages.
+Stage 1 is built for constraint acquisition, if B-Rep data is available, the openCASCADE Technology (OCCT) is used to extract constraints, otherwise, the constraint prediction from point cloud (CST-PCD) module is adopted.
+The CST-PCD is designed to rely on local features only, enabling it to generalize to unseen datasets after pre-training.
+Stage 2 leverages both point cloud and constraints for constraint feature learning, which employs both point-level and feature-level attention for effective feature extraction, enhancing the comprehension of CAD shapes.
 Finally, we built up a multi-modal classification dataset Parametric 20000, which contains B-Rep, Mesh, and point cloud.
-Given the current scarcity of labeled CAD shapes datasets, Parametric 20000 offers researchers an expanded resource for their work.
+Given the current scarcity of labeled CAD shape datasets, Parametric 20000 offers researchers an expanded resource for their work.
 Our contributions are threefold:
-  • We introduce the deep learning friendly expression of constraints.
-  • We designed the CstNet for constraint extraction and parametric feature learning.
-  • We collected and curated the Parametric 20000 dataset, which consist of mesh, point cloud, and B-Rep data.
+  • We introduced the deep learning-friendly expression of constraints.
+  • We designed the CstNet for constraint extraction and constraint feature learning.
+  • We collected and curated the Parametric 20000 dataset, which consists of mesh, point cloud, and B-Rep data.
 Our CstNet has been validated across various tasks, on the mechanical classification dataset MCB [16], our model achieved an overall accuracy improvement of 2.98 % compared to the state-of-the-art, and for mechanical segmentation dataset 360 Gallery [19], that is 6.74 %.
 2 Related Work
-For the published deep learning methods focus on CAD shape are very few, and even fewer aimed at parametric point cloud, this section will provide an overview of existing point cloud learning methods as well as CAD shapes oriented models.
-By integrating insights from above areas, a comprehensive review of the relevant works could be presented.
+For the published deep learning methods focusing on CAD shapes are very few, and even fewer aimed at parametric point clouds, this section will provide an overview of existing point cloud learning methods as well as CAD shapes-oriented models.
+By integrating insights from the above areas, a comprehensive review of the relevant works could be presented.
 2.1 Deep learning on point clouds
@@
 Besides that, it cannot be extended to 3D data segmentation or reconstruction easily.
-Voxelization-based methods convert point cloud into 3D voxels, and then processed by 3D CNNs for feature extraction [30, 47, 35, 40, 17, 20].
-However, the size and orientation of the voxels can affect computational efficiency, memory usage.
-The voxel grid quantization process may lead to geometric detail lost, and the inherent sparsity of point clouds may result in many useless voxels.
-Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [32, 33, 24, 10, 52, 2, 38, 27, 55, 56, 9, 7, 50, 49, 48, 34, 28, 26, 41, 46, 4].
+Voxelization-based methods convert point clouds into 3D voxels, and then employ 3D CNNs for feature extraction [30, 47, 35, 40, 17, 20].
+However, the size and orientation of the voxels can affect computational efficiency and memory usage.
+The voxel grid quantization process may lead to geometric detail loss, and the inherent sparsity of point clouds may result in many useless voxels.
+Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [32, 33, 24, 10, 52, 2, 38, 27, 58, 59, 9, 7, 50, 49, 48, 34, 28, 26, 41, 46, 4, 57].
 Typically, symmetric functions and k-nearest neighbors (KNN) are employed for feature extraction.
@@
 B-Rep-based methods typically involve B-Rep data conversion and feature extraction.
-Since B-Rep data cannot be directly fed into deep learning models, the challenge lies in converting B-Rep into deep learning compatible format, with graph representation is commonly used.
+Since B-Rep data cannot be directly fed into deep learning models, the challenge lies in converting B-Rep into deep learning compatible format, with graph representation commonly used.
 These approaches are commonly applied to classification [13], segmentation [19, 5], assembly [14, 43], engineering sketch generation [11], and CAD operation sequence prediction [45, 60].
@@
 These methods are primarily used for classification [29], retrieval [3], segmentation [39], and B-Rep generation [8].
-Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [8, 44, 31, 36, 25], 3D models [21, 22], or CAD operation sequences [51].
-Most of these models are generative, and utilizing transformer modules frequently.
-A key challenge is incorporating the regularity and connectivity of primitives, which typically overlooked in graphic objects analysis.
+Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [8, 44, 31, 36, 25], 3D models [21, 22, 55], or CAD operation sequences [51].
+Most of these models are generative, and frequently utilize transformer modules.
+A key challenge is incorporating the regularity and connectivity of primitives, which are typically overlooked in graphic object analysis.
 While there has been substantial progress in above domains, the engineering applications of parametric point cloud analysis demand further innovation to ensure accuracy and robustness.
@@
 3.1 Problem Statement
-Constraints is crucial for CAD shapes, can they be utilized to enhance deep learning methods’ performance?
-To answer above question, we designed the following experiments.
-The objective is to compare the prisms shown in Figure 2 with cuboids.
+Constraints are crucial for CAD shapes, can they be utilized to enhance deep learning methods’ performance?
+To answer above question, we designed the following validation experiments.
+The objective is to compare the prisms shown in Fig.
+2 with cuboids.
 In each experiment, only prisms with a specific angle are selected, creating a binary classification task between prisms and cuboids.
@@
 As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them.
-The constraint-aware model is built on PointNet++ backbone, which predict constraints and then adopt it for classification, more details in supplementary.
+The constraint-aware model is built on PointNet++ backbone, this model predicts constraints and then adopts it for classification.
+The constraint representation is defined as the relation between the point attached primitive and a reference plane, more details in supplementary.
 Figure 2: Prisms compared to cuboid.
-The experimental results are presented in Figure 3.
+The experimental results are presented in Fig.
+3.
 For prisms with angle of 50°50°50  ° and 89°89°89  °, the constraints have little improvement on the results.
 The reason lies in the 50°50°50  ° prism differs substantially from the cuboid, while the 89°89°89  ° prism is too similar to the cuboid.
-For other circumstances, incorporating constraints improves classification accuracy significantly, demonstrating that the constraint enhanced model’s comprehension on CAD shapes.
+For other circumstances, incorporating constraints improves classification accuracy significantly, demonstrating that the constraint enhanced the model’s comprehension of CAD shapes.
 Figure 3: Classification accuracy of validation experiments.
@@
 Constraint should be defined as point-wise attribute for point cloud, analogous to normal vector, as point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation.
-Following the constraint definition in CAD software is not applicable, However, once constraint is applied, the positional relations between primitives become fixed.
-Therefore, if the type of primitives and their relations are determined, the constraint on CAD shape can be defined comprehensively.
-The point-wise primitive type could be represented as the type of the primitive which the point aligned to, using one-hot encoding.
+Following the constraint definitions in CAD software is not applicable.
+However, once the constraint is applied, the positional relations between primitives become fixed.
+Therefore, if the type of primitives and their relations are determined, the constraints on CAD shapes can be defined comprehensively.
+The point-wise primitive type can be represented as the type of the primitive to which the point is attached, using one-hot encoding.
 The positional relation is defined as the combination of main axis direction and adjacency.
-Considering that the direction of primitives’ main axis is unique, as shown in Figure 4, relations such as parallel and vertical could be derived by the direction of their main axes.
+Considering that the direction of primitives’ main axis is unique, as shown in Fig.
+4, relations such as parallel and vertical could be derived from the direction of their main axes.
 However, some dimensional relations can not, such as the cylinders’ radius equivalence.
@@
 Figure 4: Main axis of primitives. Plane: normal vector. Cylinder, cone, and sphere: rotation axis.
-Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitation on the distance and size of other primitives.
-For example, in a cuboid, the adjacency and normal of the six faces ensures that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one.
-Therefore, if the adjacency combined with main axis directions, the positional relation can be determined comprehensively.
+Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitations on the distance and size of primitives.
+For example, in a cuboid, the adjacency and normal of the six faces ensure that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one.
+Therefore, if the adjacency is combined with main axis directions, the positional relation can be determined comprehensively.
 Adjacency represents the connection between primitives and cannot be expressed as point-wise attribute directly.
-The edges are the key indicator of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges.
+The edges are key indicators of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges.
 For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by two-channel one-hot vector.
-In summary, each composition of constraints in parametric point cloud is illustrated in Figure 5, and each point with constraint is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency).
+In summary, each composition of constraints in parametric point cloud is illustrated in Fig.
+5, and each point with constraint is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency).
 Figure 5: Constraint representation of parametric point cloud.
 3.3 Method Overview
-Our CstNet and module details are depicted in the top and bottom side of Figure 6 respectively.
+Our CstNet and module details are depicted in the top and bottom sides of Fig.
+6 respectively.
 It should be noted that the CST-PCD module requires pre-training, and once the pre-training is completed, its parameters are kept frozen.
-Given the locality of the constraints expression we designed, the pre-trained CST-PCD module is able to generalize to other unseen mechanical datasets directly.
+Given the locality of the constraints expression we designed, the pre-trained CST-PCD module is able to generalize to other unseen CAD shape datasets directly.
 In this work, the dataset for pre-training CST-PCD module comes from ABC [18].
@@
 Stage 1 is designed for constraint acquisition.
-When B-Rep data is available, the CST-BRep module is utilized to extract constraints; otherwise, use the CST-PCD predicts constraints.
-Stage 2 perform constraint feature learning, facilitate deeper understanding for CAD shapes.
+When B-Rep data is available, the CST-BRep module is utilized to extract constraints; otherwise, use the CST-PCD module.
+Stage 2 performs constraint feature learning, facilitating a deeper understanding of CAD shapes.
 The bottom side presents the details of module design.
@@
 For the input B-Rep data, tessellation is performed by OCCT to generate a mesh model, followed by Poisson disk sampling [23] via pymeshlab to ensure a uniform point distribution and control the point count, if point cloud generation is not needed, this step could be skipped.
-For adjacency computation, we identify valid edges firstly.
-In this study, an edge is considered invalid if the two adjacent faces are smoothly connected along the edge, this work is accomplished by OCCT.
+For adjacency computation, we identify valid edges first.
+In this study, an edge is considered invalid if the two adjacent primitives are smoothly connected along the edge, this work is accomplished by OCCT.
 Next, we compute the distance from each point to all valid edges.
 A point is marked as near an edge if the minimum distance is below a threshold (determined by the model surface area, set to $0.08\times Area/4\pi$ in this paper).
-For the primitive type and main axis direction of a point, we first determine which face it aligned with, OCCT is then used to analyze the face, providing both the face type and main axis direction.
+For the primitive type and main axis direction of a point, we first determine which primitive it attached, OCCT is then used to analyze the primitive, providing both the primitive type and main axis direction.
 In the absence of B-Rep data, such as working with mesh or voxel models, it can be converted into point cloud, after which the CST-PCD module can be used to predict constraints.
@@
 Step 1. Identify neighbor points around $\boldsymbol{p}$.
-Step 2. From all neighbor points, identify which belong to the same primitive as $\boldsymbol{p}$, these points are referred to as valid points.
-Step 3. Fit shapes such as cylinders or planes using the valid points, and determine the primitive type with the smallest fitting error.
-Step 4. Based on the the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
+Step 2. From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, these points are referred to as valid points.
+Step 3. Fit shapes such as cylinders or planes using valid points, and determine the primitive type with the smallest fitting error.
+Step 4. Based on the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
 While the above process can be carried out using traditional methods, it is highly complex, which is why a deep learning model is employed to achieve the same.
@@
 However, for constraint computation, it is more effective to search for neighbors along the shape’s surface.
-This approach captures more relevant points while reducing the influence of redundant points, as shown in Figure 7.
+This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig.
+7.
 To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary.
 Figure 7: Comparison of KNN and SurfaceKNN. Blue point: center, red point: neighbor.
-For Step 2, which involves identifying valid points from neighbors, the attention mechanism between points is employed to approximate this process, which applying different weight to neighbor points when updating the center point’s embedding.
-We designed the point attention layer based on vector self-attention, the relation between points is designed as embedding subtraction, as shown in Equation (1).
+For Step 2, which involves identifying valid points from neighbors, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding.
+We designed the point attention layer based on vector self-attention, the relation between points is designed as embedding subtraction, as shown in Eq.
+1.
 $\boldsymbol{f}^{\prime}_{i}=\sum_{f_{j}\in\mathcal{N}_{i}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{i})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{i}).$ (1)
@@
 For no global information is incorporated and most CAD shapes consist of a limited set of local geometry, such as plane, cylinder, and sphere, the CST-PCD can generalize to unseen datasets if pre-trained on a large dataset.
-As a result, our CST-PCD enables the effective use of existing unlabeled B-Rep datasets, making it possible to serve for meaningful tasks.
+As a result, our CST-PCD enables the effective use of existing unlabeled B-Rep datasets, making it possible to serve meaningful tasks.
 3.5 Constraint Feature Learning
-The stage 2 of CstNet is designed to leverage constraints to deepen the model’s understanding of CAD shapes.
-The designed architecture extract features from Primitive Type, Main Axis Direction, and Adjacency separately, attention between features is then applied to update constraint feature.
-The structure and module details are illustrated in Figure 6.
+Stage 2 of CstNet is designed to leverage constraints to deepen the model’s understanding of CAD shapes.
+The designed architecture extracts features from Primitive Type, Main Axis Direction, and Adjacency separately, attention between features is then applied to update constraint feature.
+The structure and module details are illustrated in Fig.
+6.
 The Triple MLP takes the point cloud, Primitive Type, Main Axis Direction, and Adjacency as inputs.
@@
 The outputs are the Primitive Feature, Axis Feature, and Adjacency Feature, respectively.
-In the C-MLP layer, point cloud is concatenated with Primitive Type, Main Axis Direction, Adjacency, and then feed into MLP, with the output being the Parametric Feature.
+In the C-MLP layer, point cloud is concatenated with Primitive Type, Main Axis Direction, Adjacency, and then feeds into MLP, with the output being the Parametric Feature.
 […]
-The Fea Attention layer also takes the Primitive Feature, Axis Feature, Adjacency Feature, and Parametric Feature as input, attention between features is employed to update constraint feature efficiently.
-This design is considered that different points may focus on different features, for example, points near edge may focus more on the Adjacency Feature, while others may pay more attention to the Primitive or Axis Feature.
-The Fea Attention is developed by vector attention as formulated in Equation (2).
+The Fea Attention layer also takes the Primitive Feature, Axis Feature, Adjacency Feature, and Parametric Feature as input, attention between features is employed to update the constraint feature efficiently.
+This design considers that different points may focus on different features, for example, points near edge may focus more on the Adjacency Feature, while others may pay more attention to the Primitive or Axis Feature.
+The Fea Attention is developed by vector attention as formulated in Eq.
+2.
 $\boldsymbol{f}^{\prime}_{cst}=\sum_{\boldsymbol{f}_{j}\in\mathcal{F}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{cst})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{cst}).$ (2)
@@
 4.2 Constraint Acquisition
-The extraction of constraints from B-Rep data is accomplished by CST-BRep module, with results visualized in the "GT" column of Figure 8.
-By referring the model row, this module performs well in extracting three components of constraint.
-Extract constraints from point cloud is achieved by CST-PCD module.
+The extraction of constraints from B-Rep data is accomplished by CST-BRep module, with results visualized in the "GT" column of Fig.
+8.
+By referring to the model row, this module performs well in extracting three components of constraint.
+Extracting constraints from point cloud is achieved by CST-PCD module.
 The training set is sourced from 25 trunks of ABC [18], with each trunk containing 10,000 B-Rep files.
-We utilized CST-BRep to extract constraints as Ground Truth for model training.
+We utilized CST-BRep module to extract constraints as Ground Truth for model training.
 After the training finished, the prediction accuracy of Primitive Type reached 94.23 %, for Adjacency, that is 94.62 %, and the MSE loss for Main Axis Direction is 0.0559.
-In subsequent classification and segmentation tasks, above pre-trained CST-PCD module is utilized to predict constraints with its weights frozen and no further updates.
+In subsequent classification and segmentation tasks, the above pre-trained CST-PCD module is utilized to predict constraints with its weights frozen and no further updates.
 To ensure that only point cloud is used as input across all methods in comparison, even when B-Rep data is available, the CST-BRep module will not be activated to obtain constraints.
-After CST-PCD module is pre-trained, we evaluated it on a new Trunk of ABC.
-The visualization of constraint prediction is shown in the "Pred." column of Figure 8.
+After CST-PCD module was pre-trained, we evaluated it on a new trunk of ABC.
+The visualization of constraint prediction is shown in the "Pred." column of Fig.
+8.
 After traversing this trunk, the prediction accuracy of Primitive Type reached 91.19 %, for Adjacency, that is 90.24 %, the MSE loss of Main Axis Direction is 0.0998, above demonstrating the strong capability for constraint prediction.
@@
 Main Axis Direction has been processed for visualization, details in supplementary.
-Additionally, we evaluated the pre-trained CST-PCD module on the CAD shapes dataset MCB [16] to predict constraints.
+Additionally, we evaluated the pre-trained CST-PCD module on the CAD shape dataset MCB [16] to predict constraints.
 For the MCB consists of mesh files, it was not possible to extract Ground Truth constraints for comparison.
-Based on the results visualized in Figure 9, CST-PCD demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
+Based on the results visualized in Fig.
+9, CST-PCD demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
 Figure 9: Constraint prediction on MCB.
 4.3 Classification
-The classification results on the Parametric 20000 dataset are presented in Table 1, where our CstNet outperformed other methods across both metrics.
+The classification results on the Parametric 20000 dataset are presented in Tab.
+1, where our CstNet outperformed other methods across both metrics.
 Additionally, the instance accuracy of all methods is below 90 %, with a notable performance gap between them, indicating that this dataset poses challenges and effectively differentiates between methods.
@@
 Ours 89.93 91.04 90.32 92.32
-We conducted evaluations on the CAD shapes dataset MCB [16], with results presented in Table 2.
+We conducted evaluations on the CAD shape dataset MCB [16], with results presented in Tab.
+2.
 Our CstNet achieved the highest scores across all metrics on MCB-A, but performed less effectively on MCB-B.
 This discrepancy may stem from the composition of MCB: MCB-A includes shapes from TraceParts (69.27 %), 3D Warehouse (21.70 %), and GrabCAD (9.03 %), while MCB-B is limited to shapes from 3D Warehouse and GrabCAD.
-TraceParts provides data from manufacturing companies, whereas 3D Warehouse and GrabCAD contain shapes for general usages [16].
+TraceParts provides data from manufacturing companies, whereas 3D Warehouse and GrabCAD contain shapes for general usage [16].
 As the CST-PCD module in CstNet is pre-trained on ABC, which primarily comprises mechanical components, this pre-training may lead to poor constraint prediction performance on MCB-B, thereby influencing the classification results.
@@
 We conducted experiments on the mechanical segmentation dataset 360 Gallery [19], which contains point clouds with segmentation labels based on CAD modeling operations.
-However, this labeling approach presents disadvantage for point cloud segmentation.
+However, this labeling approach presents disadvantages for point cloud segmentation.
 For example, a cube in this dataset has four faces labeled as "ExtrudeSide" and two faces labeled as "ExtrudeEnd", which poses an unrealistic task for point cloud networks.
 Additionally, the distribution of points across different labels is highly imbalanced.
-For instance, points labeled as "ExtrudeSide" accounts for 45.94 %, while those labeled as "RevolveEnd" make up only 0.12 %.
-Despite these limitations, to our knowledge, 360 Gallery is currently the only publicly available point cloud dataset for CAD shapes segmentation task.
-The segmentation results are presented in Table 3, where our method outperforms the state-of-the-art methods on most metrics.
-The results also highlight the limitations of this 360 Gallery in terms of point cloud segmentation.
+For instance, points labeled as "ExtrudeSide" account for 45.94 %, while those labeled as "RevolveEnd" make up only 0.12 %.
+Despite these limitations, to our knowledge, 360 Gallery is currently the only publicly available point cloud dataset for CAD shape segmentation tasks.
+The segmentation results are presented in Tab.
+3, where our method outperforms the state-of-the-art methods on most metrics.
+The results also highlight the limitations of 360 Gallery in terms of point cloud segmentation.
 Apart from "ExtrudeSide" and "ExtrudeEnd", none of the models were able to differentiate the other segmentation categories except DGCNN managed to predict "RevolveSide" with a low mIoU.
@@
 OA: overall accuracy %.
-mIoU: instance mIoU.
-mIoU*: class mIoU.
-ESide: extrude side mIoU, EEnd: extrude end mIoU.
-Other segmentation categories’ mIoU are all 0 except RevolveSide mIoU of DGCNN is 9.60, more datails in supplementary.
+mIoU: instance mIoU %.
+mIoU*: class mIoU %.
+ESide: Extrude Side mIoU %, EEnd: Extrude End mIoU %.
+Other segmentation categories’ mIoU are all 0 except RevolveSide mIoU of DGCNN is 9.60 %, more datails in supplementary.
 Method OA mIoU mIoU* ESide EEnd
@@
 Ours 70.76 46.65 20.51 61.29 60.65
-A portion of the segmentation results on 360 Gallery test set are shown in Table 4, where our CstNet shows segmentation results closest to the Ground Truth.
+A portion of the segmentation results on 360 Gallery test set are shown in Tab.
+4, where our CstNet shows segmentation results closest to the Ground Truth.
 PointNet’s segmentation differs from other methods, for instance, in the third and fourth rows, PointNet mistakenly segments a single primitive into two different categories.
@@
 To validate the effectiveness of our design, we conducted a series of experiments on Parametric 20000 dataset.
-For each component of constraint, we evaluated their importance by disable one at a time.
-This operation is conducted by setting it as a zero tensor before feeding into the model.
-The results are presented in Table 5, which show that all three components contribute to improve the performance of our CstNet.
+For each component of constraint, we evaluated their importance by disabling one at a time.
+This operation is performed by setting it as a zero tensor before feeding into the model.
+The results are presented in Tab.
+5, which show that all three components contribute to improving the performance of our CstNet.
 To assess the effectiveness of SurfaceKNN, we compared it with KNN.
-From Table 5, it can be concluded that in most cases, our SurfaceKNN outperforms KNN.
-However, when Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
+From Tab.
+5, it can be concluded that in most cases, our SurfaceKNN outperforms KNN.
+However, when the Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
 Table 5: Effects on Parametric 20000 dateset with varying neighbor search and constraints. Row 3 - 6: SurfaceKNN, Row 7 - 10: KNN.
@@
 5 Conclusions
-In this paper, we studied the CAD shapes and introduced the constraints expression which suitable for deep learning methods.
+In this paper, we studied the CAD shapes and introduced the constraints expression which is suitable for deep learning methods.
 Afterward, we proposed the CstNet for extracting and learning constraint features.
 Additionally, we built up a multimodal CAD shape classification dataset Parametric 20000.
-Experimental results on both public datasets and proposed dataset demonstrate that our CstNet achieves state-of-the-art performance across multiple metrics.
+Experimental results on both the public datasets and proposed dataset demonstrate that our CstNet achieves state-of-the-art performance across multiple metrics.
 Comprehensive ablation studies further validate that each component of constraint and SurfaceKNN contributes to improve the performance of CstNet.
 CAD shapes possess numerous attributes.
-In this paper, we analyze the constraints from functional perspective.
+In this paper, we analyze the constraints from a functional perspective.
 In future work, we plan to investigate additional aspects such as part design and manufacturing processes, to achieve a deeper understanding of CAD shapes.

### v3

@@
 In future work, we plan to investigate additional aspects such as part design and manufacturing processes, to achieve a deeper understanding of CAD shapes.
+Supplementary Material
+A Overview
+This supplementary material provides additional details that supplement the main paper and includes more experimental results.
+In Sec.
+B, we present the details of network architecture and data preparation in validation experiments.
+In Sec.
+D, we show more details of our constraint representation and constraint learning methods.
+In Sec.
+C, we provide the data construction and static of Parametric 20000 dataset.
+Finally, Sec.
+E offers data process for visualization and more results of experiments.
+B Validation Experiment
+B.1 Constrain Expression
+This section introduces the constraint expression in the validation experiments.
+The constraint expression in validation experiments is defined in a simple manner: for a point, its constraint is defined by the relationship between the plane it attached and the reference plane, as shown in Fig.
+10.
+This calculation process is carried out by openCASCADE Technology (OCCT).
+The constraint is represented in a one-hot encoding, where $(\mathit{1,0,0})$ indicates perpendicular, $(\mathit{0,1,0})$ indicates parallel, and $(\mathit{0,0,1})$ indicates other constraints.
+Each point in the point cloud is expressed as $(Coordinate,constraint)$.
+For example, if the coordinate of a point is $(0.1,0.2,0.3)$, and the plane it attached is perpendicular to the reference plane, it is expressed as $(0.1,0.2,0.3,\mathit{1,0,0})$, where $(0.1,0.2,0.3)$ represents the point’s coordinate, and $(\mathit{1,0,0})$ denotes the type of constraint.
+Figure 10: Constraint expression in validation experiments. The red point’s constraint is expressed as the constraint type between the magenta plane and the reference plane, using one-hot encoding.
+B.2 Dataset Generation
+This section introduces the dataset generation procedures in the validation experiments.
+The dataset generation for the validation experiments consists of the following steps:
+  1. 1.
+    Building the parametric templates for cuboids and prisms (with angles of 50°50°50  °, 60°60°60  °, 70°70°70  °, 80°80°80  °, 82°82°82  °, 85°85°85  °, 87°87°87  °, 89°89°89  °).
+Afterwards determining the dimension range for each parameter.
+The parametric templates and the dimensions are shown in Fig.
+11.
+  2. 2.
+    Assigning random values to the parameters in parametric templates. Each parametric template is instantiated into 5,000 B-Rep format CAD shapes, which are stored in STEP files [32].
+  3. 3.
+    Using OCCT to read the STEP files and convert them into meshes. Poisson disk sampling is then applied to sample points from meshes.
+  4. 4.
+    For the sampled points, OCCT is used to calculate the plane to which each point is attached.
+The constraint type for each point is then determined based on the relationship between the attached plane and the reference plane.
+  5. 5.
+    Rotating the generated point clouds randomly along the XYZ axes, within a range of [25°-25°- 25  °, 25°25°25  °].
+  6. 6.
+    Assigning the generated point clouds to each experiments.
+The validation experiments involve binary classification of cuboids and prisms with specific angles.
+The point clouds used in each of the eight experiments are as follows (80 % for training, and 20 % for testing):
+    • Experiment 1: Cuboid $\times$ 5000, Prism 50°50°50  ° $\times$ 5000
+    • Experiment 2: Cuboid $\times$ 5000, Prism 60°60°60  ° $\times$ 5000
+    • Experiment 3: Cuboid $\times$ 5000, Prism 70°70°70  ° $\times$ 5000
+    • Experiment 4: Cuboid $\times$ 5000, Prism 80°80°80  ° $\times$ 5000
+    • Experiment 5: Cuboid $\times$ 5000, Prism 82°82°82  ° $\times$ 5000
+    • Experiment 6: Cuboid $\times$ 5000, Prism 85°85°85  ° $\times$ 5000
+    • Experiment 7: Cuboid $\times$ 5000, Prism 87°87°87  ° $\times$ 5000
+    • Experiment 8: Cuboid $\times$ 5000, Prism 89°89°89  ° $\times$ 5000
+Figure 11: Parametric templates for prisms and cuboids. $a,b,c\in[0.3,2.3]$, for prisms: $b>\frac{c}{tan\alpha}$.
+B.3 Model Design
+This section introduces the structure of the constraint-aware model used in the validation experiments.
+The constraint-aware model first employs a PointNet++ [34] backbone to predict point-wise constraints.
+These constraints are then concatenated with the point coordinates and passed through another PointNet++ backbone to determine the point cloud category.
+The model’s loss function is the sum of the classification loss and the constraint loss, as shown in Fig.
+12
+Figure 12: Constraint-aware model structure.
+C Parametric 20000 Dataset
+C.1 Dataset Construction
+This section provide details of purification process and data acquisition of Parametric 20000 dataset.
+Purification: The purification process involved removing broken, duplicated, or overly complex CAD shapes requiring more than 4,000 points for point cloud representation.
+Point cloud: The point cloud data in Parametric 20000 dataset containing about 2,200 points, point constraints is also included, each point is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency), the meshes and point clouds are generated from B-Rep data.
+C.2 Dataset Statics
+This section presents the shapes count and proportion statistics for each category in the Parametric 20000 dataset. Additionally, shapes from each category are showcased.
+Parametric 20000 dataset is divided into 80 % for training and 20 % for testing.
+The class distribution is shown in Fig. 13.
+The proportion of shapes in each category are shown in Fig. 14.
+Shapes instantiated from the parametric templates designed by our team names begin with "Num".
+Example shapes from each category are shown in Fig. 15.
+Figure 13: Quantity distribution of Parametric 20000 dataset.
+Figure 14: Shapes proportion of Parametric 20000 dataset.
+Figure 15: Example shapes for each category in Parametric 20000 dataset.
+D Constraint Learning Method Details
+D.1 Defects of General Constraint Representation
+The simple constraint expression in validation experiments is insufficient for a comprehensive definition, as selecting an appropriate reference plane for complex CAD shapes is challenging.
+Moreover, traditional constraint representations suffer from non-uniqueness, as shown in Fig.
+16, defining the relations between three planes involves multiple constraint options, therefore, this approach is not well-suited for deep learning models.
+Figure 16: Non-uniqueness of traditional constraint expression. Left: Double vertical constraints, right: Vertical and parallel constraints.
+The constraint is defined as a graph representation in some circumstances, where each node represents a primitive, and the edge between nodes represent their constraints.
+However, this approach is not well-suited for point cloud analysis.
+First, in cases where only point clouds are provided as input, the number of primitives is hard to obtain.
+Second, due to the unordered nature of point cloud, it is difficult to associate points with the corresponding primitives, making label design very challenging.
+Additionally, the graph representation essentially records the constraint relationships between each pair of primitives, for the CAD shapes contain many primitives, recording the relative positions between every pair of primitives would result in an overwhelming amount of data, making the expression of these relative positions impractical.
+D.2 Constraint Representation
+This section provides the process of Main Axis Direction and example constraint representation.
+The main axis direction is processed to ensure its uniqueness, as one axis can determine two unit vectors with opposite direction.
+For a given axis, we first randomly select one of the two vectors, denoted as $v(x,y,z)$, and then process it using the following algorithm:
+Data: Unit vector $v(x,y,z)$
+Result: Unit vector after direction unification
+1 if $z<0$ then
+2       $v=-1\times v$;
+3       Return $v$;
+4
+5else if $z==0$ then
+6       if $y<0$ then
+7             $v=-1\times v$;
+8             Return $v$;
+9
+10      else if $y==0$ then
+11             if $x<0$ then
+12                   $v=-1\times v$;
+13                   Return $v$;
+14
+15
+16
+return $v$
+A point with constraint is expressed as Fig. 17.
+Figure 17: Expression of a point with constraint.
+D.3 CST-BRep Module
+This section explains why it is necessary to compute valid edges in the CST-BRep module.
+The CST-BRep module is used to compute constraints from B-Rep data, and the computation process consists of the following steps:
+Step 1: Point cloud generation (optional).
+Step 2: Calculation of valid edges and the distance from each point to these valid edges. If the minimum distance is below a threshold, the point is classified as a edge-near point.
+Step 3: Calculation of the primitives to which each point attached, followed by analysis to determine point’s Primitive Type and the Main Axis Direction.
+In Step 2, valid edges are computed, as this helps to reduce the potential for data to mislead the deep learning model.
+The definition of valid edge is illustrated in Fig.
+18.
+This definition is based on the observation that a complete cylinder in the B-Rep data stored in STEP files is often split into two half-cylinder faces.
+As a result, two generatrices form edges on the cylindrical surface, but the deep learning model cannot predict these edges, as shown in Fig.
+19.
+Furthermore, in other cases, edges corresponding to smoothly connected faces are also difficult to identify, leading to confusion for deep learning models.
+Figure 18: Zebra stripes of primitives corresponding to invalid and valid edges. Left: G1 or higher continuous at the invalid edges, right: G0 continuous at the valid edges.
+Figure 19: Cylinder in B-Rep representation.
+A complete cylindrical surface is represented as a combination of two half-cylinder surface, resulting in edges along the generatrices that cannot be recognized, thereby causing confusion for deep learning model.
+D.4 CST-PCD Module
+This section provides a detailed explanation of the valid points in the computational steps that inspire the design of the CST-PCD Module.
+The CST-PCD Module is used to predict constraints from point clouds. The design of the CST-PCD Module is inspired by the following calculation process:
+For a point $\boldsymbol{p}$ in point cloud, its constraint could be calculated by the following steps:
+  1. 1.
+    Identify neighbor points around $\boldsymbol{p}$.
+  2. 2.
+    From all neighbor points, identify those that belong to the same primitive as $\boldsymbol{p}$; these points are referred to as valid points.
+  3. 3.
+    Fit shapes such as cylinders or planes using valid points, and determine the primitive type with the smallest fitting error.
+  4. 4.
+    Based on the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
+In these steps, it is necessary to compute the valid points.
+We define the valid point as the neighbors that lie on the same primitive as the center point $\boldsymbol{p}$, as visualized in Fig.
+20.
+These valid points can be used to determine whether the center point is near an edge, as well as to calculate the main axis direction and primitive type of the attached primitive, while the invalid points do not contribute to these calculations.
+Figure 20: Valid and invalid points for constraint calculation.
+D.5 SurfaceKNN
+This section introduces the algorithm of SurfaceKNN.
+SurfaceKNN based on the assumption that when KNN is applied with a small number of neighbors, the output points are searched along the shape’s surface.
+This assumption has been validated in most cases.
+Leveraging this assumption, SurfaceKNN is accomplished by applying KNN with small number of neighbors iteratively.
+During each iteration, KNN is used to search a small set of neighboring points, which then serve as new central points for subsequent searches.
+This process continues until the desired number of points is obtained, as illustrated in Fig.
+21.
+Figure 21: Algorithm of SurfaceKNN.
+E Classification Segmentation Experiments
+E.1 Process to Visualize Main Axis Direction
+The main axis direction of a point, denoted as $v=(x,y,z)$, is visualized as $R=\frac{x+1}{2}$, $G=\frac{y+1}{2}$, and $B=\frac{z+1}{2}$.
+Since the main axis direction is represented as a unit vector, $-1\leqslant x,y,z\leqslant 1$, which results in $0\leqslant\frac{x+1}{2},\frac{y+1}{2},\frac{z+1}{2}\leqslant 1$.
+E.2 Constraint Visualization of CST-BRep
+More constraints extracted by the CST-BRep module from the B-Rep are visualized in Fig. 22.
+Figure 22: More constraint extraction results from the B-Rep data. ADJ: Adjacency, PT: Primitive Type, MAD: Main Axis Direction.
+E.3 Segmentation
+The full segmentation results of investigated methods on 360 gallery are shown in Tab. 6.
+Table 6: Part Segmentation Results on 360 Gallery.
+OA: overall accuracy %.
+mIoU: instance mIoU %.
+mIoU*: class mIoU %.
+Segmentation categories from left to right are: Extrude Side, Extrude End, Cut Side, Cut End, Fillet, Chamfer, Revolve Side, Revolve End.
+PN: PointNet [33].
+PN2: PointNet++ [34].
+DGN: DGCNN [42].
+3DN: 3DGCN [26].
+Method PN PN2 DGN 3DN Ours
+OA 42.13 58.18 64.02 60.58 70.76
+mIoU 18.89 26.98 35.54 33.03 46.65
+mIoU* 7.04 11.27 14.36 12.23 20.51
+[Uncaptioned image] 19.83 49.80 54.48 50.67 61.29
+[Uncaptioned image] 36.47 40.36 50.82 47.20 60.65
+[Uncaptioned image] 0 0 0 0 0
+[Uncaptioned image] 0 0 0 0 0
+[Uncaptioned image] 0 0 0 0 0
+[Uncaptioned image] 0 0 0 0 0
+[Uncaptioned image] 0 0 9.60 0 0
+[Uncaptioned image] 0 0 0 0 0

### v5

@@
-Constraint Learning for Parametric Point Cloud
+Constraint-Aware Feature Learning for Parametric Point Cloud
 Abstract
-Parametric point clouds are sampled from CAD shapes, and have become increasingly prevalent in industrial manufacturing.
-However, most existing point cloud learning methods focus on the geometric features, such as developing efficient convolution operations, overlooking the important attribute of constraints inherent in CAD shapes, which limits these methods’ ability to comprehend CAD shapes fully.
-To address this issue, we analyzed the effect of constraints, and proposed its deep learning-friendly representation, after that, the Constraint Feature Learning Network (CstNet) was developed to extract and leverage constraints.
-Our CstNet includes two stages.
-Stage 1 extracts constraints from B-Rep data or point cloud.
-Stage 2 leverages coordinates and constraints to enhance the comprehension of CAD shapes.
-Additionally, we built up the Parametric 20,000 Multi-modal Dataset for the scarcity of labeled B-Rep datasets.
-Experiments demonstrate that our CstNet achieved state-of-the-art performance on both public and proposed CAD shape datasets.
-To the best of our knowledge, CstNet is the first constraint-based learning method tailored for CAD shape analysis.
-Project page: https://cstnetwork.github.io/
+Parametric point clouds are sampled from CAD shapes and are becoming increasingly common in industrial manufacturing.
+Most existing CAD-specific deep learning methods only focus on geometric features, while overlooking constraints which are inherent and important in CAD shapes.
+This limits their ability to discern CAD shapes with similar appearance but different constraints.
+To tackle this challenge, we first analyze the constraint importance via a simple validation experiment.
+Then, we introduce a deep learning-friendly constraints representation with three vectorized components, and design a constraint-aware feature learning network (CstNet), which includes two stages.
+Stage 1 extracts constraint feature from B-Rep data or point cloud based on shape local information.
+It enables better generalization ability to unseen dataset after model pre-training.
+Stage 2 employs attention layers to adaptively adjust the weights of three constraints’ components.
+It facilitates the effective utilization of constraints.
+In addition, we built the first multi-modal parametric-purpose dataset, i.e.
+Param20K, comprising about 20K shape instances of 75 classes.
+On this dataset, we performed the classification and rotation robustness experiments, and CstNet achieved 3.52% and 26.17% absolute improvements in instance accuracy over the state-of-the-art methods, respectively.
+To the best of our knowledge, CstNet is the first constraint-aware deep learning method tailored for parametric point cloud analysis in CAD domain.
+Our project page with source code is available at: https://cstnetwork.github.io/.
 1 Introduction
-Parametric point clouds are derived from parametric templates, which consist of primitives, constraints, and dimensional parameters.
-By specifying all parameters, the templates can be instantiated into CAD shapes, as illustrated in Fig.
-1.
-Parametric point clouds provide detailed information that is essential for various engineering applications, such as model design and product machining.
-In recent years, deep learning methods have emerged as powerful tools for point cloud analysis, with models such as PointNet [33], PointNet++ [34], and PointCNN [24] demonstrating significant success across various tasks.
-These models are primarily designed and evaluated on graphic objects, such as animals, trees, and other free-form shapes, where shape comprehension relies on geometric features visible in contours.
-However, some CAD shapes, despite exhibiting minimal visual differences, serve distinct purposes due to variations in their functional regions.
-This similarity in appearance poses challenges in distinguishing them based on geometric features.
-Notably, variations in the functional regions often introduce differing constraints, as these regions need to fit closely with other parts to perform functions such as transmission and positioning.
-Thus, distinguishing them from a constraint-based perspective becomes more feasible, as illustrated in Fig.
-1.
 Figure 1: Motivation.
-Some similar CAD shapes serve different functions.
-The eccentric wheel has a rotation axis that is offset from its outer cylindrical surface, facilitating precise positioning or tension adjustment.
-In contrast, the flywheel has a coaxial rotation axis with its outer cylindrical surface and is used to store rotational kinetic energy.
-Although these shapes are challenging to distinguish visually, they can be easily differentiated from the perspective of constraints.
-Additionally, the limited availability of B-Rep datasets poses barriers to deep learning on CAD shapes.
-Most existing CAD shape datasets consist of mesh files, such as MCB [16] and ESB [12].
-While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information.
-In contrast, B-Rep data [19, 57] serves as the native representation of CAD shapes and is therefore more suitable for dataset construction.
-However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files.
-Although the ABC [18] contains a large number of B-Rep files, it remains unlabeled.
-To facilitate the constraint feature learning of CAD shapes, we conducted studies on deep learning methods and dataset development.
-Traditional constraint definitions are not well-suited for deep learning, as they vary across constraint types and involve multiple possible combinations to represent primitives’ relations.
-To address this challenge, we proposed a novel constraint representation as point-wise local attributes.
-Afterward, we developed the CstNet for constraint extraction and feature learning, which consist of two stages.
-Stage 1 is built for constraint acquisition, if B-Rep data is available, the openCASCADE Technology (OCCT) is used to extract constraints, otherwise, the constraint prediction from point cloud (CST-PCD) module is adopted.
-The CST-PCD is designed to rely on local features only, enabling it to generalize to unseen datasets after pre-training.
-Stage 2 leverages both point cloud and constraints for constraint feature learning, which employs both point-level and feature-level attention for effective feature extraction, enhancing the comprehension of CAD shapes.
-Finally, we built up a multi-modal classification dataset Parametric 20000, which contains B-Rep, Mesh, and point cloud.
-Given the current scarcity of labeled CAD shape datasets, Parametric 20000 offers researchers an expanded resource for their work.
-Our contributions are threefold:
-  • We introduced the deep learning-friendly expression of constraints.
-  • We designed the CstNet for constraint extraction and constraint feature learning.
-  • We collected and curated the Parametric 20000 dataset, which consists of mesh, point cloud, and B-Rep data.
-Our CstNet has been validated across various tasks, on the mechanical classification dataset MCB [16], our model achieved an overall accuracy improvement of 2.98 % compared to the state-of-the-art, and for mechanical segmentation dataset 360 Gallery [19], that is 6.74 %.
+Some CAD shapes are visually similar but serve different functions.
+The eccentric wheel has a distance constraint, i.e.
+its rotation axis is offset from outer cylindrical surface.
+It is used for precise positioning or tension adjustment.
+In contrast, the flywheel has a coaxial constraint, i.e.
+its rotation axis is coaxial with outer cylindrical surface.
+It is used to store rotational kinetic energy.
+While these shapes are challenging to distinguish visually, they can be effectively differentiated from constraints.
+Parametric point clouds are sampled from CAD shape instances or captured from various industrial scenarios by common digital devices.
+As illustrated in Fig.
+1, CAD shape instances are typically instantiated from parametric templates, which comprise primitives, constraints, and shape parameters.
+By specifying a series of parameter configurations, a template can be instantiated into a family of CAD shape instances [62].
+In industrial domain, many CAD shapes are visual similar but have different constraints in their functional regions.
+This poses challenges for existing point cloud deep learning methods which based on geometric features only.
+Thus, a constraint-aware feature learning method for parametric point cloud is fundamental for a wide range of industrial applications, such as engineering model design and product manufacturing.
+First, most existing studies [58, 41, 26, 65] focus on primitive regularity (e.g.
+uniform appearance and distinct boundaries).
+They usually leverage geometric features only, but ignore those constraints between primitives.
+Although these methods have been effectively applied in point cloud segmentation and primitive reconstruction, they inherently limit their discrimination capability for CAD shapes that are visually similar yet functionally distinct.
+While some studies discussed constraints in 2D engineering sketches [8, 40], the considered primitives in 2D sketches are supposed available and generally few in number.
+Thus, these approaches can be hardly extended to 3D parametric point clouds.
+Second, most CAD-specific methods rely on specialized datasets with additional labels (e.g.
+primitive, parameter, and normal labels).
+However, point clouds are typically acquired via LiDAR or RGB-D cameras, and in most cases only contain coordinate information.
+It thus restricts the practical applicability of these approaches.
+Third, there are still numerous challenges in constraint-aware point cloud analysis.
+Constraints represent the spatial relationships between primitives, but the number and types of primitives cannot be directly obtained from point clouds.
+Predicting primitives from point clouds also poses challenges.
+The unordered nature of point clouds results in unordered outcomes, leading to a mismatch between the predicted primitives and those in the constraint labels.
+In addition, processing relations between every pair of primitives would produce an overwhelming amount of data, since some CAD shapes contain many primitives.
+Assuming that the constraints have been obtained, effectively leveraging them remains a significant challenge.
+To address the aforementioned issues, we transform the traditional constraint representation (e.g.
+constraints in SolidWorks or UG [3]) into three vectorized sub-features: primitive’s Main Axis Direction (MAD), Adjacency (Adj), and Primitive Type (PT).
+Given the completeness of traditional constraint representation, these three sub-features can effectively capture constraints.
+Representing constraints with MAD, Adj, PT converts the relations between primitive pairs into relations between primitives and a common reference, thereby significantly reducing the total number of relations.
+Moreover, the MAD, Adj, PT can be inferred by local point patches, and the local features of CAD shapes are predominantly composed of primitives such as planes and cylinders.
+Consequently, a model that relies solely on local features can generalize to unseen datasets after pre-trained, this property eliminates the reliance on constraint labels during practical applications.
+To mitigate the impact of point cloud unordered nature, we encode MAD, Adj, PT as point-wise features.
+As a result, each point with constraints in point cloud is represented as (X, Y, Z, MAD, Adj, PT).
+The aforementioned processing transform the constraints into deep learning-friendly form.
+After that, we designed the constraint prediction from point cloud (CstPnt) module to obtain MAD, Adj, PT.
+The CstPnt module is based solely on local features, experiments show that it exhibits strong generalization ability on unseen datasets, and its training data can be easily obtained by processing unlabeled B-Rep data.
+Following constraint acquisition, we separately extract features from MAD, Adj, PT, and apply attention mechanism to weight these features, thereby facilitating their effective utilization.
+We quantitatively evaluate CstNet on ABC [20] and Param20K dataset: constraints prediction, classification, and rotation robustness evaluation.
+For constraints prediction, our CstPnt module (in CstNet Stage 1) surpasses the state-of-the-art (SOTA) by around 5% and 14% instance accuracy in ABC (test set) and Param20K (unseen dataset), respectively.
+For classification and rotation robustness on Param20K, CstNet outperforms the SOTA by 3.52% and 26.17% instance accuracy, respectively.
+In summary, our contributions are threefold:
+  • We introduced a deep learning-friendly vectorized representation and designed the first constraint-aware feature learning network CstNet for effective constraint acquisition and leveraging for parametric point cloud.
+  • We built the first parametric-purpose Param20K dataset, which consists of about 20K shape instances of 75 classes.
+Each instance include mesh, point cloud, and B-Rep data.
+Given the current scarcity of labeled CAD shape datasets, Param20K offers a solid basis for future studies in parametric CAD domains.
+  • We conducted extensive experiments on constraint acquisition, classification, robustness evaluation and ablation studies.
+The performance of CstNet has a significant improvement compared with SOTA methods which usually focus on geometric data only.
 2 Related Work
-For the published deep learning methods focusing on CAD shapes are very few, and even fewer aimed at parametric point clouds, this section will provide an overview of existing point cloud learning methods as well as CAD shapes-oriented models.
-By integrating insights from the above areas, a comprehensive review of the relevant works could be presented.
-2.1 Deep learning on point clouds
-Point cloud deep learning can be categorized into three types based on data processing approaches: view-based, voxel-based, and point-based methods.
+General point cloud analysis.
 View-based methods typically project the point cloud from multiple views, the obtained figures are then fed in convolutional neural networks (CNNs) for feature extraction [42, 17, 7, 60, 59, 47].
-However, these methods face significant challenges when applied to large-scale scanned data, primarily due to the difficulty of covering entire scenes from single-point perspectives, and the choice of projection views significantly impacts recognition performance.
-Besides that, it cannot be extended to 3D data segmentation or reconstruction easily.
+How to address information loss during the projection and optimizing viewpoints play a critical role.
 Voxelization-based methods convert point clouds into 3D voxels, and then employ 3D CNNs for feature extraction [33, 52, 39, 45, 19, 22].
-However, the size and orientation of the voxels can affect computational efficiency and memory usage.
-The voxel grid quantization process may lead to geometric detail loss, and the inherent sparsity of point clouds may result in many useless voxels.
+The size and orientation of 3D voxels significantly influences the accuracy of these methods, and information loss occurs during the conversion to voxels.
 Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [36, 37, 27, 12, 57, 2, 43, 30, 64, 66, 11, 9, 55, 54, 53, 38, 31, 29, 46, 51, 5, 63].
-Typically, symmetric functions and k-nearest neighbors (KNN) are employed for feature extraction.
-However, the unordered nature and transformation invariance of point clouds pose challenges for efficient feature extraction.
-2.2 Deep learning on CAD shapes
-CAD shapes deep learning can be categorized into three types based on the input data format: B-Rep-based learning, mesh-based learning, and sketch-based learning.
-B-Rep-based methods typically involve B-Rep data conversion and feature extraction.
-Since B-Rep data cannot be directly fed into deep learning models, the challenge lies in converting B-Rep into deep learning compatible format, with graph representation commonly used.
+The unordered nature and rotational invariance of point clouds make it challenging to employ methods similar to traditional CNNs for feature extraction, and kNN is a commonly adopted approach.
+Parametric point cloud analysis.
+The most extensive application of these methods lies in segmentation and primitive fitting [58, 41, 26, 65].
+CAD shapes are typically composed of regular primitives, such as plane, cylinder, cone, and the variety of such elements is limited.
+Therefore, leveraging primitives’ distinct shape features enables these methods to achieve superior segmentation and primitive fitting outcomes compared to general approaches.
+These methods generally predicting primitives parameters, and subsequently fed into downstream modules, these modules may incorporate specific primitive fitting algorithms.
+Despite their superior ability to exploit the shape features of individual primitives, these methods overlooked the constraints between primitives.
+This oversight limited their ability to distinguish CAD shapes that share similar appearances but serve different functions.
+Moreover, the primitive parameters predicting procedure necessitates the corresponding parameter labels.
+In practical, when sampling point clouds from CAD shapes, only coordinates can be obtained in most cases, which limits the applicability of these methods.
+Other CAD-specific studies.
+B-Rep-based methods typically involve B-Rep data conversion and feature extraction, the challenge lies in converting B-Rep into deep learning compatible format.
 These approaches are commonly applied to classification [15], segmentation [21, 6], assembly [16, 48], engineering sketch generation [13], and CAD operation sequence prediction [50, 67].
-Mesh-based learning methods use triangle models as input, such as STL and OBJ.
-Compared to B-Rep, mesh representations lack boundary details, which prevents these methods from leveraging data conversion for improved performance.
-As a result, contributions often rely on mechanical datasets provided by the authors.
+Mesh-based learning methods use triangle models as input, such as STL and OBJ, contributions often rely on mechanical datasets provided by the authors.
 These methods are primarily used for classification [32], retrieval [4], segmentation [44], and B-Rep generation [10].
 Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [10, 49, 34, 40, 28], 3D models [23, 24, 62], or CAD operation sequences [56].
-Most of these models are generative, and frequently utilize transformer modules.
-A key challenge is incorporating the regularity and connectivity of primitives, which are typically overlooked in graphic object analysis.
-While there has been substantial progress in above domains, the engineering applications of parametric point cloud analysis demand further innovation to ensure accuracy and robustness.
+A key challenge is incorporating the regularity and connectivity of primitives, which is typically overlooked in graphic object analysis.
+While there has been substantial progress in CAD-specific studies, further innovation on constraint learning is demanded to ensure higher accuracy and robustness.
 3 Method
-3.1 Problem Statement
-Constraints are crucial for CAD shapes, can they be utilized to enhance deep learning methods’ performance?
-To answer above question, we designed the following validation experiments.
+3.1 Is Constraint Important?
+We aim to utilize constraints to distinguish visually similar but functionally distinct CAD shapes.
+To answer whether constraints could enhance deep learning methods’ performance, we designed the following validation experiments.
 The objective is to compare the prisms shown in Fig.
@@
 As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them.
-The constraint-aware model is built on PointNet++ backbone, this model predicts constraints and then adopts it for classification.
-The constraint representation is defined as the relation between the point attached primitive and a reference plane, more details in supplementary.
+The constraint-aware model is built on PointNet++ backbone, which predicts constraints and then adopts it for classification.
+The constraints are represented as the relation between the point attached primitive and a reference plane, more details in supplementary.
 Figure 2: Prisms compared to cuboid.
-The experimental results are presented in Fig.
-3.
-For prisms with angle of 50°50°50  ° and 89°89°89  °, the constraints have little improvement on the results.
-The reason lies in the 50°50°50  ° prism differs substantially from the cuboid, while the 89°89°89  ° prism is too similar to the cuboid.
-For other circumstances, incorporating constraints improves classification accuracy significantly, demonstrating that the constraint enhanced the model’s comprehension of CAD shapes.
+From results presented in Fig.
+3, for prisms with angles from 50°50°50  ° to 87°87°87  °, incorporating constraints improves classification accuracy significantly, while the 89°89°89  ° prisms are too similar to the cuboids, all methods cannot distinguish them.
 Figure 3: Classification accuracy of validation experiments.
@@
 3.2 Constraint Representation
-Constraint should be defined as point-wise attribute for point cloud, analogous to normal vector, as point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation.
+In general, constraints should be defined as point-wise attribute for point clouds, analogous to normal vector.
+As point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation.
 Following the constraint definitions in CAD software is not applicable.
@@
 Thus, additional data is required to supplement this information.
-Figure 4: Main axis of primitives. Plane: normal vector. Cylinder, cone, and sphere: rotation axis.
 Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitations on the distance and size of primitives.
@@
 In summary, each composition of constraints in parametric point cloud is illustrated in Fig.
-5, and each point with constraint is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency).
+5, and each point with constraint is expressed as (X, Y, Z, MAD, Adj, PT).
+Figure 4: Main Axis Direction of primitives. Plane: normal vector. Cylinder, cone, sphere: rotation axis. Main Axis Direction requires unique process, details are provided in Supplementary.
+We represent constraints by MAD, Adj, PT, is this approach sufficient to handle most constraints?
+For PT, commercial CAD systems typically classify primitives into plan, cylinder, cone, sphere, and free-form surface.
+Among these, plane, cylinder, and cone are primarily influenced by constraints, while spherical and free-form surfaces rely heavily on their geometric features.
+Therefore, considering plane, cylinder, and cone is sufficient for constraint representation.
+For sphere and free-form surface, the Stage 2 of CstNet also incorporate geometric features (feed in coordinates).
+MAD correspond to directional constraints (e.g.
+parallel and vertical), and MAD combined with PT represent dimensional constraints (e.g.
+distance and tangent).
+Through MAD and Adj, most common constraint types can be covered.
+Therefore, the combination of MAD, Adj, and PT is sufficient to represent the majority of constraints.
 Figure 5: Constraint representation of parametric point cloud.
-3.3 Method Overview
-Our CstNet and module details are depicted in the top and bottom sides of Fig.
-6 respectively.
-It should be noted that the CST-PCD module requires pre-training, and once the pre-training is completed, its parameters are kept frozen.
-Given the locality of the constraints expression we designed, the pre-trained CST-PCD module is able to generalize to other unseen CAD shape datasets directly.
-In this work, the dataset for pre-training CST-PCD module comes from ABC [18].
+3.3 Constraint Acquisition
+The Stage 1 of CstNet is designed to obtain constraints, as shown in the left of Fig.
+7.
+When B-Rep data is available, the constraint computation from B-Rep (CstBRep) module is developed to obtain constraints.
+For the input B-Rep data, tessellation is performed by OCCT to generate a mesh model, followed by Poisson disk sampling [25] via pymeshlab to ensure a uniform point distribution and control the point count, if point cloud generation is not needed, this step could be skipped.
+For adjacency computation, we identify valid edges firstly.
+An edge is considered invalid if the two adjacent primitives are smoothly connected along the edge.
+Searching valid edges is accomplished by OCCT.
+Next, we compute the distance from each point to all valid edges.
+A point is marked as near an edge if the minimum distance is below a threshold (set to $0.08\times ShapeArea/4\pi$ in this paper).
+For the primitive type and main axis direction of a point, we first determine which primitive it attached, OCCT is then used to analyze the primitive, providing both the primitive type and main axis direction.
+In the absence of B-Rep data, such as working with mesh or voxel models, it can be converted to point cloud, after which the CstPnt module can be used to predict constraints.
+Since each point’s constraints are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary.
+The constraint computation of a point $\boldsymbol{p}$ could be divided into the following steps: Step 1: Identify neighbor points around $\boldsymbol{p}$.
+Step 2: From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, store in $\boldsymbol{\mathrm{P_{valid}}}$.
+Step 3: Fit shapes such as cylinders or planes using $\boldsymbol{\mathrm{P_{valid}}}$, and determine the primitive type with the smallest fitting error.
+Step 4: Based on $\boldsymbol{\mathrm{P_{valid}}}$ and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
+While the above process can be carried out using traditional methods, it is highly complex, which is why a deep learning model is employed to achieve the same.
+The design of CstPnt module is inspired by the above steps, with details in Fig.
+7.
+For Step 1, the traditional approach is to use the k-Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere.
+However, for constraint computation, it is more effective to search for neighbors along the shape’s surface.
+This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig.
+6.
+To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary.
+For Step 2, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding, the relation between points is designed as embedding subtraction, as shown in Eq.
+1.
+Step 3 and Step 4 involve computing MAD and PT, both steps can be accomplished by MLPs.
+We employed a U-Net-like structure for better data utilization.
+$\boldsymbol{f}^{\prime}_{i}=\sum_{f_{j}\in\mathcal{N}_{i}}\rho\Bigl{(}\rm{MLP}\bigl{(}\mathrm{Q}(\boldsymbol{f}_{\mathit{i}})-\mathrm{K}(\boldsymbol{f}_{\mathit{j}})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{\mathit{i}}).$ (1)
+Where $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication.
+$\mathrm{Q}(\boldsymbol{f})$, $\mathrm{K}(\boldsymbol{f})$, and $\mathrm{V}(\boldsymbol{f})$ are defined in similar structure by concatenating $\boldsymbol{f}$ with the positional encoding $\Delta$, and then feeding it into MLPs, where $\Delta$ is defined as $\Delta=\boldsymbol{p}_{i}-\boldsymbol{p}_{j}$.
+Figure 6: Comparison of KNN and SurfaceKNN. Blue point: center of neighbor search, red point: searched neighbors.
+The structure of CstPnt reveals that only local information is utilized, which aligns with the local nature of the constraint representation we designed.
+For no global information is incorporated and most CAD shapes consist of a limited set of primitives, such as plane, cylinder, and sphere, the CstPnt could generalize to unseen datasets if pre-trained on a large scale dataset.
+The training data is derived from unlabeled B-Rep data, as a result, our CstPnt enables the effective use of existing large scale unlabeled B-Rep datasets, making it possible to serve meaningful tasks.
 Figure 7: Overview of CstNet.
-The top side illustrates the overall architecture of CstNet, comprising two stages.
+The top side illustrates the overall architecture of CstNet, which comprising two stages.
 Stage 1 is designed for constraint acquisition.
-When B-Rep data is available, the CST-BRep module is utilized to extract constraints; otherwise, use the CST-PCD module.
+When B-Rep data is available, the CstBRep module is utilized to extract constraints; otherwise, use the CstPnt module.
 Stage 2 performs constraint feature learning, facilitating a deeper understanding of CAD shapes.
 The bottom side presents the details of module design.
-3.4 Constraint Acquisition
-When B-Rep data is available, the CST-BRep module can be used to compute constraints.
-For the input B-Rep data, tessellation is performed by OCCT to generate a mesh model, followed by Poisson disk sampling [23] via pymeshlab to ensure a uniform point distribution and control the point count, if point cloud generation is not needed, this step could be skipped.
-For adjacency computation, we identify valid edges first.
-In this study, an edge is considered invalid if the two adjacent primitives are smoothly connected along the edge, this work is accomplished by OCCT.
-Next, we compute the distance from each point to all valid edges.
-A point is marked as near an edge if the minimum distance is below a threshold (determined by the model surface area, set to $0.08\times Area/4\pi$ in this paper).
-For the primitive type and main axis direction of a point, we first determine which primitive it attached, OCCT is then used to analyze the primitive, providing both the primitive type and main axis direction.
-In the absence of B-Rep data, such as working with mesh or voxel models, it can be converted into point cloud, after which the CST-PCD module can be used to predict constraints.
-Since each point’s constraints are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary.
-The constraint computation of a point $\boldsymbol{p}$ could be divided into the following steps:
-Step 1. Identify neighbor points around $\boldsymbol{p}$.
-Step 2. From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, these points are referred to as valid points.
-Step 3. Fit shapes such as cylinders or planes using valid points, and determine the primitive type with the smallest fitting error.
-Step 4. Based on the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
-While the above process can be carried out using traditional methods, it is highly complex, which is why a deep learning model is employed to achieve the same.
-The design of CST-PCD module is inspired by the aforementioned calculation process.
-For Step 1, which involves finding neighbor points, the traditional approach is to use the K Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere.
-However, for constraint computation, it is more effective to search for neighbors along the shape’s surface.
-This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig.
+In the CstPnt module, the final MLP consists of three parallel MLPs with the same input, where the outputs correspond to Main Axis Direction, Adjacency, and Primitive Type.
+3.4 Constraint-Aware Feature Learning
+The Stage 2 of CstNet is designed to leverage constraints effectively, which extracts features from MAD, Adj, PT separately, and attention between features is then applied to update constraint feature, with details in Fig.
 7.
-To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary.
-Figure 7: Comparison of KNN and SurfaceKNN. Blue point: center, red point: neighbor.
-For Step 2, which involves identifying valid points from neighbors, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding.
-We designed the point attention layer based on vector self-attention, the relation between points is designed as embedding subtraction, as shown in Eq.
-1.
-$\boldsymbol{f}^{\prime}_{i}=\sum_{f_{j}\in\mathcal{N}_{i}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{i})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{i}).$ (1)
-Where $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication.
-$\mathrm{Q}(\boldsymbol{f})$, $\mathrm{K}(\boldsymbol{f})$, and $\mathrm{V}(\boldsymbol{f})$ are defined in the same way by concatenating $\boldsymbol{f}$ with the positional encoding $\Delta$ and then feeding it into MLP, where $\Delta$ is defined as $\Delta=\boldsymbol{p}_{i}-\boldsymbol{p}_{j}$.
-Step 3 and Step 4 involve determining the Primitive Type and Main Axis Direction, both steps can be accomplished by MLPs.
-We employed a U-Net-like structure to enhance information utilization.
-In the CST-PCD module, C-MLP refers to concatenating the input tensors and feeding them into MLPs.
-The final MLP in this module consists of three parallel MLPs with the same input, where the outputs correspond to Primitive Type, Main Axis Direction, and Adjacency.
-The structure of CST-PCD module reveals that only local information is utilized, which aligns with the inherently local nature of the constraint representation we designed.
-For no global information is incorporated and most CAD shapes consist of a limited set of local geometry, such as plane, cylinder, and sphere, the CST-PCD can generalize to unseen datasets if pre-trained on a large dataset.
-As a result, our CST-PCD enables the effective use of existing unlabeled B-Rep datasets, making it possible to serve meaningful tasks.
-3.5 Constraint Feature Learning
-Stage 2 of CstNet is designed to leverage constraints to deepen the model’s understanding of CAD shapes.
-The designed architecture extracts features from Primitive Type, Main Axis Direction, and Adjacency separately, attention between features is then applied to update constraint feature.
-The structure and module details are illustrated in Fig.
-6.
-The Triple MLP takes the point cloud, Primitive Type, Main Axis Direction, and Adjacency as inputs.
-In this module, the point cloud is concatenated with the Primitive Type, Main Axis Direction, and Adjacency separately before being fed into three individual MLPs.
-The outputs are the Primitive Feature, Axis Feature, and Adjacency Feature, respectively.
-In the C-MLP layer, point cloud is concatenated with Primitive Type, Main Axis Direction, Adjacency, and then feeds into MLP, with the output being the Parametric Feature.
 […]
-The Fea Attention layer also takes the Primitive Feature, Axis Feature, Adjacency Feature, and Parametric Feature as input, attention between features is employed to update the constraint feature efficiently.
-This design considers that different points may focus on different features, for example, points near edge may focus more on the Adjacency Feature, while others may pay more attention to the Primitive or Axis Feature.
-The Fea Attention is developed by vector attention as formulated in Eq.
-2.
-$\boldsymbol{f}^{\prime}_{cst}=\sum_{\boldsymbol{f}_{j}\in\mathcal{F}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{cst})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{cst}).$ (2)
-In which =$\mathcal{F}=\{\boldsymbol{f}_{pmt},\boldsymbol{f}_{axs},\boldsymbol{f}_{adj},\boldsymbol{f}_{cst}\}$, $\boldsymbol{f}_{pmt}$ is Primitive Feature, $\boldsymbol{f}_{axs}$ is Axis Feature, $\boldsymbol{f}_{adj}$ is Adjacency Feature, and $\boldsymbol{f}_{cst}$ is Constraint Feature.
-$\Delta$ is defined as one-hot encoding of $\boldsymbol{f}_{j}$ type.
-For downstream tasks, if classification is required, the final constraint feature can be fed into MLP and softmax.
-For segmentation tasks, upsampling and skip link concatenation similar to those in PointNet++ [34] can be applied.
+$\boldsymbol{f}^{\prime}_{\rm{cst}}=\sum_{\boldsymbol{f}_{j}\in\mathcal{F}}\rho\Bigl{(}\rm{MLP}\bigl{(}\mathrm{Q}(\boldsymbol{f}_{\rm{cst}})-\mathrm{K}(\boldsymbol{f}_{\mathit{j}})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{\rm{cst}}).$ (2)
+In which =$\mathcal{F}=\{\boldsymbol{f}_{\rm{mad}},\boldsymbol{f}_{\rm{adj}},\boldsymbol{f}_{\rm{pt}},\boldsymbol{f}_{\rm{cst}}\}$, corresponding to Axis Feature, Adjacency Feature, Primitive Feature, and Constraint Feature.
+$\Delta$ is the one-hot encoding of $\boldsymbol{f}_{j}$ type.
+For downstream classification tasks, the constraint features obtained in each module will be concatenated by residual connection and feed into maxpoling, MLPs and softmax.
+Constraint features obtained by different modules may differ in features’ number and channel, based on the coordinates of the final constraint features, we identify the correspondence from different modules for further processing.
 4 Experiments
 4.1 Experimental Settings
-In general, we train CstNet using the Negative Log Likelihood Loss (NLLLoss) and optimized by Adam optimization for 200 epochs, with a weight decay of 0.0001 and an initial learning rate of 0.0001.
+We train CstNet using the Negative Log Likelihood Loss and optimized by Adam optimization for 200 epochs, with a weight decay of 0.0001 and an initial learning rate of 0.0001.
 We use StepLR learning rate scheduler, with step size of 20 and gamma of 0.7.
@@
 All experiments are performed on GeForce RTX 4090 GPU.
-Parametric 20000 dataset: our Parametric 20000 dataset contains 19,739 multi-modal instances categorized into 75 classes.
+Param20K dataset Param20K contains 19,739 multi-modal instances categorized into 75 classes.
 Each instance includes B-Rep data, mesh, and point cloud.
@@
 More details in the supplementary.
-Sources 1 (36.05 %): Downloaded and purified from TraceParts [16]. The CAD shapes on TraceParts are collected from manufacturing companies.
-Sources 2 (24.12 %): Provided by a CAD company [^1] https://www.newdimchina.com/ that collaborates with our team. These shapes are all mechanical components.
-Sources 3 (39.83 %): Instantiated from parametric templates designed by our team. These templates are designed based on common mechanical structures.
-Our Parametric 20000 dataset encompasses various common formats of CAD shapes, making it suitable for a wide range of tasks.
+Sources 1 (36.05 %): Downloaded and purified from TraceParts [18].
+The CAD shapes on TraceParts are collected from manufacturing companies.
+Sources 2 (24.12 %): Provided by a CAD company [^1] https://www.newdimchina.com/ that collaborates with our team.
+These shapes are all mechanical components.
+Sources 3 (39.83 %): Instantiated from parametric templates designed by our team.
+These templates are designed based on common mechanical structures.
+Our Param20K dataset encompasses various common formats of CAD shapes, making it suitable for a wide range of tasks.
 The careful selection and purification process undertaken by our team further ensures its high quality.
 4.2 Constraint Acquisition
-The extraction of constraints from B-Rep data is accomplished by CST-BRep module, with results visualized in the "GT" column of Fig.
+The extraction of constraints from B-Rep data is accomplished by the CstBRep module, with results visualized in the Ground Truth column of Fig.
 8.
-By referring to the model row, this module performs well in extracting three components of constraint.
-Extracting constraints from point cloud is achieved by CST-PCD module.
+By referring to the model row, CstBRep performs well in extracting MAD, Adj, and PT from B-Rep data.
+Predicting MAD, Adj, and PT from point cloud is achieved by the CstPnt module.
 The training set is sourced from 25 trunks of ABC [20], with each trunk containing 10,000 B-Rep files.
-We utilized CST-BRep module to extract constraints as Ground Truth for model training.
-After the training finished, the prediction accuracy of Primitive Type reached 94.23 %, for Adjacency, that is 94.62 %, and the MSE loss for Main Axis Direction is 0.0559.
-In subsequent classification and segmentation tasks, the above pre-trained CST-PCD module is utilized to predict constraints with its weights frozen and no further updates.
-To ensure that only point cloud is used as input across all methods in comparison, even when B-Rep data is available, the CST-BRep module will not be activated to obtain constraints.
-After CST-PCD module was pre-trained, we evaluated it on a new trunk of ABC.
-The visualization of constraint prediction is shown in the "Pred." column of Fig.
-8.
-After traversing this trunk, the prediction accuracy of Primitive Type reached 91.19 %, for Adjacency, that is 90.24 %, the MSE loss of Main Axis Direction is 0.0998, above demonstrating the strong capability for constraint prediction.
-Figure 8: Constraint prediction on ABC.
-MAD: Main Axis Direction.
-PT: Primitive Type.
-ADJ: Adjacency (whether point is near edge).
-Main Axis Direction has been processed for visualization, details in supplementary.
-Additionally, we evaluated the pre-trained CST-PCD module on the CAD shape dataset MCB [16] to predict constraints.
-For the MCB consists of mesh files, it was not possible to extract Ground Truth constraints for comparison.
-Based on the results visualized in Fig.
-9, CST-PCD demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
-Figure 9: Constraint prediction on MCB.
+We utilized CstBRep module to extract constraints as Ground Truth for model training.
+After the training finished, the MSE loss of MAD is 0.0559, the prediction accuracy of Adj and PT reached 94.62% and 94.23%, respectively.
+In subsequent experiments, the above pre-trained CstPnt module is utilized to predict MAD, Adj, and PT with its weights frozen and no further updates.
+To ensure that only point cloud is used as input across all methods when comparing, even when B-Rep data is available, the CstBRep module will not be activated to obtain MAD, Adj, and PT.
+After CstPnt was pre-trained, we evaluated it on a ABC new trunk and Param20K dataset.
+The ABC new trunk was treated as the test set, while the Param20K was considered as an unseen CAD shapes dataset, more results on other public available CAD shapes datasets in supplementary.
+Since there is no existing model that directly predicts MAD, Adj, and PT, we modified the published CAD-specific methods [58, 41] by adding MLPs as prediction head to enable these predictions.
+The experiment results are shown in Tab.
+1, from which our CstPnt achieves the highest accuracy in predicting MAD, Adj, and PT.
+Furthermore, from ABC to Param20K, all methods’ performance decreased, but our method dropped least accuracy.
+ParSeNet and HPNet may be negatively affected by global features, which impaired their performance on unseen datasets.
+Fig.
+8 presents the visualization of constraint prediction results, which shown that our method is closest to the ground truth.
+Furthermore, from ABC to Paeam20K, our method demonstrates excellent generalization performance on unseen datasets.
+In contrast, while ParSeNet and HPNet exhibit acceptable prediction results on the ABC new trunk, they performed unsatisfactory on the unseen Param20K dataset.
+Table 1: Constraint prediction on ABC (new trunk) and Param20K. MAD: MSE loss of Main Axis Direction. Adj, PT: prediction accuracy of Adjacency and Primitive Type.
+[…]
+ABC ParSeNet [41] 0.1356 84.89 85.22
+HPNet [58] 0.1576 82.63 79.15
+Ours 0.0998 90.24 91.19
+Param20K ParSeNet [41] 0.2247 81.42 59.75
+HPNet [58] 0.2570 78.60 57.66
+Ours 0.1390 87.95 86.52
+Figure 8: Constraint prediction visualization on ABC and Param20K. MAD is processed for visualization, details in supplementary.
 4.3 Classification
-The classification results on the Parametric 20000 dataset are presented in Tab.
-1, where our CstNet outperformed other methods across both metrics.
-Additionally, the instance accuracy of all methods is below 90 %, with a notable performance gap between them, indicating that this dataset poses challenges and effectively differentiates between methods.
-Table 1: Classification results on Parametric 20000. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
+To validate the ability of our CstNet in identifying CAD shapes that are visual similar but different in constraints, we conducted experiments on the Prism-Cuboid datasets presented in the Problem Statement section Sec.
+3.1.
+Results shown in Fig.
+9, in which our CstNet achieved best instance accuracy in distinguish 50°50°50  ° to 87°87°87  ° prisms and cuboids, which demonstrate the strong ability of CstNet in recognizing shapes with similar appearance but constraints different.
+For the binary classification of 89°89°89  ° prisms and cuboids, all methods achieved approximately 50% instance accuracy, this is attributed to the differences between them are so subtle that no one is capable to distinguishing them.
+Figure 9: Classification results on Prism-Cuboid dataset.
+The classification of prisms and cuboids is predominantly influenced by constraints, while the comprehension of common CAD shapes necessitates the integration of both constraint and geometric features.
+CstNet achieves this by concatenating XYZ with constraint in Stage 2, thereby integrating geometric features into constraint features.
+To evaluate CstNet’s performance on common CAD shapes, we conducted experiments on Param20K dataset (more results on public CAD shapes dataset in Supplementary), results are presented in Tab.
+2, where our CstNet outperformed other methods across both metrics, validated the effectiveness of utilizing constraint and geometric features.
+Table 2: Classification results on Param20K dataset. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
 Method Acc Acc* F1 mAP
-PointCNN [24] 61.92 55.29 53.97 40.99
-PointNet [33] 75.61 74.08 74.23 82.15
-PointNet++ [34] 82.89 86.80 82.31 88.05
-PointConv [47] 80.53 82.10 81.86 86.54
-DGCNN [42] 85.99 88.20 87.20 89.20
-3DGCN [26] 86.19 86.31 85.37 87.64
-Ours 89.93 91.04 90.32 92.32
-We conducted evaluations on the CAD shape dataset MCB [16], with results presented in Tab.
-2.
-Our CstNet achieved the highest scores across all metrics on MCB-A, but performed less effectively on MCB-B.
-This discrepancy may stem from the composition of MCB: MCB-A includes shapes from TraceParts (69.27 %), 3D Warehouse (21.70 %), and GrabCAD (9.03 %), while MCB-B is limited to shapes from 3D Warehouse and GrabCAD.
-TraceParts provides data from manufacturing companies, whereas 3D Warehouse and GrabCAD contain shapes for general usage [16].
-As the CST-PCD module in CstNet is pre-trained on ABC, which primarily comprises mechanical components, this pre-training may lead to poor constraint prediction performance on MCB-B, thereby influencing the classification results.
-Table 2: Classification results on MCB. A / B column: results of MCB A / B subset.
-Method Acc.instance % Acc.class % F1-score mAP %
-A B A B A B A B
-PointCNN [24] 93.89 93.67 81.85 86.80 83.86 88.63 90.13 93.86
-PointNet [33] 86.78 80.36 67.70 65.30 86.55 79.89 74.08 71.66
-PointNet++ [34] 87.45 93.91 73.68 87.97 88.32 73.45 91.33 91.37
-SpiderCNN [53] 93.59 89.31 79.70 79.29 81.30 80.72 86.64 82.47
-PointConv [47] 93.25 90.64 80.24 80.17 71.31 76.07 82.19 86.09
-DGCNN [42] 92.54 90.75 74.47 78.17 76.12 79.69 74.27 79.73
-3DGCN [26] 93.71 92.13 78.71 84.31 84.59 79.24 84.35 84.35
-Ours 96.87 92.61 89.21 85.94 89.85 85.68 93.17 91.25
-4.4 Segmentation
-We conducted experiments on the mechanical segmentation dataset 360 Gallery [19], which contains point clouds with segmentation labels based on CAD modeling operations.
-However, this labeling approach presents disadvantages for point cloud segmentation.
-For example, a cube in this dataset has four faces labeled as "ExtrudeSide" and two faces labeled as "ExtrudeEnd", which poses an unrealistic task for point cloud networks.
-Additionally, the distribution of points across different labels is highly imbalanced.
-For instance, points labeled as "ExtrudeSide" account for 45.94 %, while those labeled as "RevolveEnd" make up only 0.12 %.
-Despite these limitations, to our knowledge, 360 Gallery is currently the only publicly available point cloud dataset for CAD shape segmentation tasks.
-The segmentation results are presented in Tab.
-3, where our method outperforms the state-of-the-art methods on most metrics.
-The results also highlight the limitations of 360 Gallery in terms of point cloud segmentation.
-Apart from "ExtrudeSide" and "ExtrudeEnd", none of the models were able to differentiate the other segmentation categories except DGCNN managed to predict "RevolveSide" with a low mIoU.
-Table 3: Part Segmentation Results on 360 Gallery.
-PN: PointNet.
-PN2: PointNet++.
-DGN: DGCNN.
-3DN: 3DGCN.
-OA: overall accuracy %.
-mIoU: instance mIoU %.
-mIoU*: class mIoU %.
-ESide: Extrude Side mIoU %, EEnd: Extrude End mIoU %.
-Other segmentation categories’ mIoU are all 0 except RevolveSide mIoU of DGCNN is 9.60 %, more datails in supplementary.
-Method OA mIoU mIoU* ESide EEnd
-PN [33] 42.13 18.89 7.04 19.83 36.47
-PN2 [34] 58.18 26.98 11.27 49.80 40.36
-DGN [42] 64.02 35.54 14.36 54.48 50.82
-3DN [26] 60.58 33.03 12.23 50.67 47.20
-Ours 70.76 46.65 20.51 61.29 60.65
-A portion of the segmentation results on 360 Gallery test set are shown in Tab.
-4, where our CstNet shows segmentation results closest to the Ground Truth.
-PointNet’s segmentation differs from other methods, for instance, in the third and fourth rows, PointNet mistakenly segments a single primitive into two different categories.
-This discrepancy likely arises from the model’s inability to effectively capture local features.
-Table 4: Visualization of part segmentation on 360 Gallery. GT: ground truth. PN: PointNet [33]. PN2: PointNet++ [34]. DGN: DGCNN [42]. 3DN: 3DGCN [26].
-GT PN PN2 DGN 3DN Ours
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+PointCNN [27] 80.25 76.03 74.64 83.38
+PointNet [36] 81.30 83.21 82.06 85.18
+PointNet++ [37] 83.70 86.37 85.30 87.94
+PointConv [51] 82.30 85.62 84.91 87.92
+DGCNN [46] 85.40 87.28 86.43 89.17
+3DGCN [29] 86.42 88.25 87.63 89.81
+Ours 89.94 91.06 90.34 92.72
+To access the robustness of CstNet, we evaluated it under rotated point clouds.
+Since point clouds are pre-processed by aligning their centroids to the origin and scaling their ranges into a unit cube, the performance of all methods is unaffected by point cloud shifts and scaling.
+Comparison results depicted in Fig.
+10, from which our CstNet demonstrates strongest robustness to point cloud rotation.
+The dashed line in Fig.
+10 show that the performance of CstPnt module remains nearly unaffected by rotation, owing to it learned various local features from ABC dataset.
+Figure 10: Classification results on rotated Param20K.
+The training set remained unchanged and the test set was rotated by a certain angle along the +Z direction.
+Solid line: left y axis, dashed line: right y axis.
+Acc.MAD = 1 - MAD.MSELoss.
 4.4 Ablation Studies
-Our model leverages all three components of constraint: Primitive Type, Main Axis Direction, and Adjacency.
-Additionally, we utilized the SurfaceKNN to obtain more suitable neighbors for each point.
-To validate the effectiveness of our design, we conducted a series of experiments on Parametric 20000 dataset.
-For each component of constraint, we evaluated their importance by disabling one at a time.
-This operation is performed by setting it as a zero tensor before feeding into the model.
-The results are presented in Tab.
-5, which show that all three components contribute to improving the performance of our CstNet.
+Our CstNet leverages all three components of constraint: MAD, Adj, and PT.
+Additionally, we utilized the SurfaceKNN to obtain more efficient neighbors for each point.
+To validate the effectiveness of our design, we conducted a series of ablation studies on Param20K dataset.
+For each component of constraint, we evaluated their importance by disabling one or all at a time.
+This operation is performed by setting the disabled components as a zero tensor before feeding into CstNet.
+Classification results are presented in Tab.
+3, which show that all three components contribute to improving the performance of our CstNet.
 To assess the effectiveness of SurfaceKNN, we compared it with KNN.
 From Tab.
-5, it can be concluded that in most cases, our SurfaceKNN outperforms KNN.
+3, it can be concluded that in most cases, our SurfaceKNN outperforms KNN.
 However, when the Primitive Type is disabled, the results are similar between the two methods, indicating that SurfaceKNN primarily enhances the effectiveness of the Primitive Feature.
-Table 5: Effects on Parametric 20000 dateset with varying neighbor search and constraints. Row 3 - 6: SurfaceKNN, Row 7 - 10: KNN.
+Table 3: Effects on varying neighbor search and constraint components. Upper half: SurfaceKNN, lower half: KNN.
 Constraint Acc Acc* F1 mAP
 MAD Adj PT
-89.93 91.04 90.32 92.32
+89.94 91.06 90.34 92.72
 86.32 89.05 88.38 90.59
@@
 88.68 88.94 89.01 91.22
+83.99 84.05 82.54 86.15
 89.03 90.00 89.22 91.43
@@
 88.73 88.97 88.91 91.21
+83.19 82.54 81.46 86.34
+To validate the effectiveness of Stage 2 backbone, we conducted a comprehensive evaluation by feeding (x, y, z, MAD, Adj, PT) into multiple classification backbone for comparison.
+The results are presented in Tab.
+4, from which our method demonstrates best performance under both predicted and label constraints, indicating that our design achieves the best utilization of constraints.
+Furthermore, although the accuracy using predicted constraints is lower than that of using label constraints, it still shows a significant improvement compared to the baselines without constraints.
+This highlights the effectiveness of the predicted constraints in improving model performance.
+Table 4: Effects on Stage2 Backbone. Upper half: predicted constraints from CstPnt, lower: label constraints.
+Backbone-S2 Acc Acc* F1 mAP
+PointNet [36] 83.83 83.64 83.20 87.27
+PointNet++ [37] 85.52 88.11 87.18 90.79
+DGCNN [46] 86.81 88.70 87.70 91.35
+3DGCN [29] 87.46 89.59 87.97 91.92
+Ours 89.94 91.06 90.34 92.72
+PointNet [36] 86.24 85.99 86.33 91.00
+PointNet++ [37] 86.60 88.82 87.71 90.74
+DGCNN [46] 87.72 89.14 88.78 91.16
+3DGCN [29] 89.29 89.81 89.00 92.62
+Ours 91.80 92.52 91.76 94.20
 5 Conclusions
-In this paper, we studied the CAD shapes and introduced the constraints expression which is suitable for deep learning methods.
-Afterward, we proposed the CstNet for extracting and learning constraint features.
-Additionally, we built up a multimodal CAD shape classification dataset Parametric 20000.
-Experimental results on both the public datasets and proposed dataset demonstrate that our CstNet achieves state-of-the-art performance across multiple metrics.
-Comprehensive ablation studies further validate that each component of constraint and SurfaceKNN contributes to improve the performance of CstNet.
+In this paper, we studied the constraints of CAD shapes and introduced its deep learning-friendly expression.
+Afterward, we proposed the CstNet for extracting and leveraging constraints.
+Additionally, we built up a multi-modal CAD shape dataset Param20K.
+Extensive experiment results demonstrate that our CstNet achieves state-of-the-art performance on both accuracy and robustness.
+Comprehensive ablation studies further validate the effectiveness of constraints, SurfaceKNN, and network backbone.
 CAD shapes possess numerous attributes.
 In this paper, we analyze the constraints from a functional perspective.
-In future work, we plan to investigate additional aspects such as part design and manufacturing processes, to achieve a deeper understanding of CAD shapes.
+In future work, we plan to investigate additional aspects such as part design and manufacturing procedure, to achieve a deeper understanding of CAD shapes.
 Supplementary Material
-A Overview
+F Overview
 This supplementary material provides additional details that supplement the main paper and includes more experimental results.
 In Sec.
-B, we present the details of network architecture and data preparation in validation experiments.
+G, we present the details of network architecture and data preparation in validation experiments.
 In Sec.
-D, we show more details of our constraint representation and constraint learning methods.
+H, we provide the data construction and static of Param20K dataset.
+Finally, Sec.
+I offers data process for visualization and more results of constraint acquisition and classification experiments.
 In Sec.
-C, we provide the data construction and static of Parametric 20000 dataset.
-Finally, Sec.
-E offers data process for visualization and more results of experiments.
-B Validation Experiment
-B.1 Constrain Expression
+J, we show more details of our constraint representation and constraint learning methods.
+G Validation Experiment
+G.1 Constraint Expression
 This section introduces the constraint expression in the validation experiments.
 The constraint expression in validation experiments is defined in a simple manner: for a point, its constraint is defined by the relationship between the plane it attached and the reference plane, as shown in Fig.
-10.
+11.
 This calculation process is carried out by openCASCADE Technology (OCCT).
@@
 Figure 11: Constraint expression in validation experiments. The red point’s constraint is expressed as the constraint type between the magenta plane and the reference plane, using one-hot encoding.
-B.2 Dataset Generation
+G.2 Dataset Generation
 This section introduces the dataset generation procedures in the validation experiments.
@@
 The parametric templates and the dimensions are shown in Fig.
-11.
+12.
   2. 2.
@@
 Figure 12: Parametric templates for prisms and cuboids. $a,b,c\in[0.3,2.3]$, for prisms: $b>\frac{c}{tan\alpha}$.
-B.3 Model Design
+G.3 Model Design
 This section introduces the structure of the constraint-aware model used in the validation experiments.
@@
 The model’s loss function is the sum of the classification loss and the constraint loss, as shown in Fig.
-12
+13
 Figure 13: Constraint-aware model structure.
-C Parametric 20000 Dataset
-C.1 Dataset Construction
-This section provide details of purification process and data acquisition of Parametric 20000 dataset.
+H Param20K Dataset
+H.1 Why build up the Param20K dataset?
+The limited availability of B-Rep datasets poses barriers to deep learning on CAD shapes.
+Most existing CAD shape datasets consist of mesh files, such as MCB [18] and ESB [14].
+While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information.
+In contrast, B-Rep data [21, 61] serves as the native representation of CAD shapes and is therefore more suitable for dataset construction.
+However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files.
+Although the ABC [20] contains a large number of B-Rep files, it remains unlabeled.
+H.2 Dataset Construction
+This section provide details of purification process and data acquisition of Param20K dataset.
 Purification: The purification process involved removing broken, duplicated, or overly complex CAD shapes requiring more than 4,000 points for point cloud representation.
-Point cloud: The point cloud data in Parametric 20000 dataset containing about 2,200 points, point constraints is also included, each point is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency), the meshes and point clouds are generated from B-Rep data.
-C.2 Dataset Statics
-This section presents the shapes count and proportion statistics for each category in the Parametric 20000 dataset. Additionally, shapes from each category are showcased.
-Parametric 20000 dataset is divided into 80 % for training and 20 % for testing.
+Point cloud: The point cloud data in Param20K dataset containing about 2,200 points, point constraints is also included, each point is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency), the meshes and point clouds are generated from B-Rep data.
+H.3 Dataset Statics
+This section presents the shapes count and proportion statistics for each category in the Param20K dataset. Additionally, shapes from each category are showcased.
+Param20K dataset is divided into 80 % for training and 20 % for testing.
 The class distribution is shown in Fig. 14.
 The proportion of shapes in each category are shown in Fig. 15.
-Shapes instantiated from the parametric templates designed by our team names begin with "Num".
+Shapes instantiated from the parametric templates designed by our team names begin with ”Num”.
 Example shapes from each category are shown in Fig. 16.
-Figure 13: Quantity distribution of Parametric 20000 dataset.
-Figure 14: Shapes proportion of Parametric 20000 dataset.
-Figure 15: Example shapes for each category in Parametric 20000 dataset.
-D Constraint Learning Method Details
-D.1 Defects of General Constraint Representation
+Figure 14: Quantity distribution of Param20K dataset.
+Figure 15: Shapes proportion of Param20K dataset.
+Figure 16: Example shapes for each category in Param20K dataset.
+I Constraint Acquisition and Classification Experiments
+I.1 Process to Visualize Main Axis Direction
+The main axis direction of a point, denoted as $v=(x,y,z)$, is visualized as $R=\frac{x+1}{2}$, $G=\frac{y+1}{2}$, and $B=\frac{z+1}{2}$.
+Since the main axis direction is represented as a unit vector, $-1\leqslant x,y,z\leqslant 1$, which results in $0\leqslant\frac{x+1}{2},\frac{y+1}{2},\frac{z+1}{2}\leqslant 1$.
+I.2 Constraint Visualization of CstBRep
+More constraints extracted by the CstBRep module from the B-Rep are visualized in Fig. 17.
+Figure 17: More constraint extraction results from the B-Rep data. ADJ: Adjacency, PMT: Primitive Type, MAD: Main Axis Direction.
+I.3 Constraint acquisition on MCB
+We evaluated the pre-trained CstPnt module and ParSeNet [41], HPNet [58] on the CAD shape dataset MCB [18] to predict constraints.
+For the MCB consists of mesh files, it was not possible to extract Ground Truth constraints for comparison, therefore no ground truth presented in Fig.
+18, and no numerical results are available.
+Based on the results visualized in Fig.
+18, CstPnt demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
+Figure 18: Constraint prediction on MCB.
+I.4 Classification on MCB
+We conducted evaluations on the CAD shape dataset MCB [18], with results presented in Tab.
+5, from which our CstNet achieved the highest scores across all metrics, conclusions similar with experiments on Param20K dataset.
+Table 5: Classification results on MCB. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
+Method Acc Acc* F1 mAP
+PointCNN [27] 93.89 81.85 83.86 90.13
+PointNet [36] 86.78 67.70 86.55 74.08
+PointNet++ [37] 87.45 73.68 88.32 91.33
+SpiderCNN [57] 93.59 79.70 81.30 86.64
+PointConv [51] 93.25 80.24 71.31 82.19
+DGCNN [46] 92.54 74.47 76.12 74.27
+3DGCN [29] 93.71 78.71 84.59 84.35
+Ours 96.87 89.21 89.85 93.17
+J Constraint Learning Method Details
+J.1 Defects of General Constraint Representation
 The simple constraint expression in validation experiments is insufficient for a comprehensive definition, as selecting an appropriate reference plane for complex CAD shapes is challenging.
 Moreover, traditional constraint representations suffer from non-uniqueness, as shown in Fig.
-16, defining the relations between three planes involves multiple constraint options, therefore, this approach is not well-suited for deep learning models.
+19, defining the relations between three planes involves multiple constraint options, therefore, this approach is not well-suited for deep learning models.
 Figure 19: Non-uniqueness of traditional constraint expression. Left: Double vertical constraints, right: Vertical and parallel constraints.
@@
 Additionally, the graph representation essentially records the constraint relationships between each pair of primitives, for the CAD shapes contain many primitives, recording the relative positions between every pair of primitives would result in an overwhelming amount of data, making the expression of these relative positions impractical.
-D.2 Constraint Representation
+J.2 Unique Process of MAD
 This section provides the process of Main Axis Direction and example constraint representation.
@@
 Figure 20: Expression of a point with constraint.
-D.3 CST-BRep Module
-This section explains why it is necessary to compute valid edges in the CST-BRep module.
-The CST-BRep module is used to compute constraints from B-Rep data, and the computation process consists of the following steps:
+J.3 CstBRep Module
+This section explains why it is necessary to compute valid edges in the CstBRep module.
+The CstBRep module is used to compute constraints from B-Rep data, and the computation process consists of the following steps:
 Step 1: Point cloud generation (optional).
@@
 The definition of valid edge is illustrated in Fig.
-18.
+21.
 This definition is based on the observation that a complete cylinder in the B-Rep data stored in STEP files is often split into two half-cylinder faces.
 As a result, two generatrices form edges on the cylindrical surface, but the deep learning model cannot predict these edges, as shown in Fig.
-19.
+22.
 Furthermore, in other cases, edges corresponding to smoothly connected faces are also difficult to identify, leading to confusion for deep learning models.
@@
 A complete cylindrical surface is represented as a combination of two half-cylinder surface, resulting in edges along the generatrices that cannot be recognized, thereby causing confusion for deep learning model.
-D.4 CST-PCD Module
-This section provides a detailed explanation of the valid points in the computational steps that inspire the design of the CST-PCD Module.
-The CST-PCD Module is used to predict constraints from point clouds. The design of the CST-PCD Module is inspired by the following calculation process:
+J.4 CstPnt Module
+This section provides a detailed explanation of the valid points in the computational steps that inspire the design of the CstPnt Module.
+The CstPnt Module is used to predict constraints from point clouds. The design of the CstPnt Module is inspired by the following calculation process:
 For a point $\boldsymbol{p}$ in point cloud, its constraint could be calculated by the following steps:
@@
 We define the valid point as the neighbors that lie on the same primitive as the center point $\boldsymbol{p}$, as visualized in Fig.
-20.
+23.
 These valid points can be used to determine whether the center point is near an edge, as well as to calculate the main axis direction and primitive type of the attached primitive, while the invalid points do not contribute to these calculations.
 Figure 23: Valid and invalid points for constraint calculation.
-D.5 SurfaceKNN
+J.5 SurfaceKNN
 This section introduces the algorithm of SurfaceKNN.
@@
 This process continues until the desired number of points is obtained, as illustrated in Fig.
-21.
+24.
 Figure 24: Algorithm of SurfaceKNN.
-E Classification Segmentation Experiments
-E.1 Process to Visualize Main Axis Direction
-The main axis direction of a point, denoted as $v=(x,y,z)$, is visualized as $R=\frac{x+1}{2}$, $G=\frac{y+1}{2}$, and $B=\frac{z+1}{2}$.
-Since the main axis direction is represented as a unit vector, $-1\leqslant x,y,z\leqslant 1$, which results in $0\leqslant\frac{x+1}{2},\frac{y+1}{2},\frac{z+1}{2}\leqslant 1$.
-E.2 Constraint Visualization of CST-BRep
-More constraints extracted by the CST-BRep module from the B-Rep are visualized in Fig. 22.
-Figure 22: More constraint extraction results from the B-Rep data. ADJ: Adjacency, PT: Primitive Type, MAD: Main Axis Direction.
-E.3 Segmentation
-The full segmentation results of investigated methods on 360 gallery are shown in Tab. 6.
-Table 6: Part Segmentation Results on 360 Gallery.
-OA: overall accuracy %.
-mIoU: instance mIoU %.
-mIoU*: class mIoU %.
-Segmentation categories from left to right are: Extrude Side, Extrude End, Cut Side, Cut End, Fillet, Chamfer, Revolve Side, Revolve End.
-PN: PointNet [33].
-PN2: PointNet++ [34].
-DGN: DGCNN [42].
-3DN: 3DGCN [26].
-Method PN PN2 DGN 3DN Ours
-OA 42.13 58.18 64.02 60.58 70.76
-mIoU 18.89 26.98 35.54 33.03 46.65
-mIoU* 7.04 11.27 14.36 12.23 20.51
-[Uncaptioned image] 19.83 49.80 54.48 50.67 61.29
-[Uncaptioned image] 36.47 40.36 50.82 47.20 60.65
-[Uncaptioned image] 0 0 0 0 0
-[Uncaptioned image] 0 0 0 0 0
-[Uncaptioned image] 0 0 0 0 0
-[Uncaptioned image] 0 0 0 0 0
-[Uncaptioned image] 0 0 9.60 0 0
-[Uncaptioned image] 0 0 0 0 0

### v6

@@
 Parametric point clouds are sampled from CAD shapes and are becoming increasingly common in industrial manufacturing.
-Most existing CAD-specific deep learning methods only focus on geometric features, while overlooking constraints which are inherent and important in CAD shapes.
-This limits their ability to discern CAD shapes with similar appearance but different constraints.
-To tackle this challenge, we first analyze the constraint importance via a simple validation experiment.
-Then, we introduce a deep learning-friendly constraints representation with three vectorized components, and design a constraint-aware feature learning network (CstNet), which includes two stages.
-Stage 1 extracts constraint feature from B-Rep data or point cloud based on shape local information.
-It enables better generalization ability to unseen dataset after model pre-training.
+Most CAD-specific deep learning methods focus on geometric features, while overlooking constraints inherent in CAD shapes.
+This limits their ability to discern CAD shapes with similar appearances but different constraints.
+To tackle this challenge, we first analyze the constraint importance via simple validation experiments.
+Then, we introduce a deep learning-friendly constraints representation with three components, and design a constraint-aware feature learning network (CstNet), which includes two stages.
+Stage 1 extracts constraint representation from BRep data or point cloud based on local features.
+It enables better generalization ability to unseen dataset after pre-training.
 Stage 2 employs attention layers to adaptively adjust the weights of three constraints’ components.
@@
 In addition, we built the first multi-modal parametric-purpose dataset, i.e.
-Param20K, comprising about 20K shape instances of 75 classes.
-On this dataset, we performed the classification and rotation robustness experiments, and CstNet achieved 3.52% and 26.17% absolute improvements in instance accuracy over the state-of-the-art methods, respectively.
-To the best of our knowledge, CstNet is the first constraint-aware deep learning method tailored for parametric point cloud analysis in CAD domain.
+Param20K, comprising about 20K CAD instances of 75 classes.
+On this dataset, CstNet achieved 3.49% (classification) and 26.17% (rotation robustness) accuracy improvements over the state-of-the-art.
+To the best of our knowledge, CstNet is the first constraint-aware deep learning method tailored for parametric point cloud analysis.
 Our project page with source code is available at: https://cstnetwork.github.io/.
+[…]
 1 Introduction
@@
 While these shapes are challenging to distinguish visually, they can be effectively differentiated from constraints.
-Parametric point clouds are sampled from CAD shape instances or captured from various industrial scenarios by common digital devices.
+Parametric point clouds are sampled from CAD shapes or captured from various industrial scenarios by common digital devices.
 As illustrated in Fig.
-1, CAD shape instances are typically instantiated from parametric templates, which comprise primitives, constraints, and shape parameters.
-By specifying a series of parameter configurations, a template can be instantiated into a family of CAD shape instances [62].
-In industrial domain, many CAD shapes are visual similar but have different constraints in their functional regions.
-This poses challenges for existing point cloud deep learning methods which based on geometric features only.
-Thus, a constraint-aware feature learning method for parametric point cloud is fundamental for a wide range of industrial applications, such as engineering model design and product manufacturing.
+1, CAD shapes are typically instantiated from parametric templates, which comprise primitives, constraints, and shape parameters.
+By specifying a series of parameter configurations, a template can be instantiated into a family of CAD shapes [67].
+In industrial domain, many CAD shapes are visual similar but have different constraints in their functional regions, examples shown in Fig.
+1.
+This poses challenges for geometric features-based point cloud deep learning methods.
+Thus, a constraint-aware feature learning method for parametric point cloud is fundamental for a wide range of industrial applications, such as engineering model design [67, 18].
 First, most existing studies [63, 43, 28, 70] focus on primitive regularity (e.g.
 uniform appearance and distinct boundaries).
-They usually leverage geometric features only, but ignore those constraints between primitives.
+They usually leverage geometric features only, but ignore constraints between primitives.
 Although these methods have been effectively applied in point cloud segmentation and primitive reconstruction, they inherently limit their discrimination capability for CAD shapes that are visually similar yet functionally distinct.
@@
 Constraints represent the spatial relationships between primitives, but the number and types of primitives cannot be directly obtained from point clouds.
-Predicting primitives from point clouds also poses challenges.
-The unordered nature of point clouds results in unordered outcomes, leading to a mismatch between the predicted primitives and those in the constraint labels.
-In addition, processing relations between every pair of primitives would produce an overwhelming amount of data, since some CAD shapes contain many primitives.
+Predicting primitives from point clouds also poses challenges, the point cloud unordered nature results in unordered outcomes, leading to a mismatch between the predictions and labels.
+In addition, processing relations between every primitive pairs would produce an overwhelming data volume, since some CAD shapes contain many primitives.
 Assuming that the constraints have been obtained, effectively leveraging them remains a significant challenge.
 To address the aforementioned issues, we transform the traditional constraint representation (e.g.
-constraints in SolidWorks or UG [3]) into three vectorized sub-features: primitive’s Main Axis Direction (MAD), Adjacency (Adj), and Primitive Type (PT).
-Given the completeness of traditional constraint representation, these three sub-features can effectively capture constraints.
-Representing constraints with MAD, Adj, PT converts the relations between primitive pairs into relations between primitives and a common reference, thereby significantly reducing the total number of relations.
-Moreover, the MAD, Adj, PT can be inferred by local point patches, and the local features of CAD shapes are predominantly composed of primitives such as planes and cylinders.
+constraints in SolidWorks or UG [4]) into three components: primitive’s Main Axis Direction (MAD), Adjacency (Adj), and Primitive Type (PT).
+Representing constraints with MAD-Adj-PT converts the relations between primitive pairs into relations between primitives and a common reference, thereby significantly reducing the data volume.
+Moreover, MAD-Adj-PT can be inferred by local point patches, and the local features of CAD shapes are predominantly composed of primitives such as planes and cylinders.
 Consequently, a model that relies solely on local features can generalize to unseen datasets after pre-trained, this property eliminates the reliance on constraint labels during practical applications.
-To mitigate the impact of point cloud unordered nature, we encode MAD, Adj, PT as point-wise features.
-As a result, each point with constraints in point cloud is represented as (X, Y, Z, MAD, Adj, PT).
-The aforementioned processing transform the constraints into deep learning-friendly form.
-After that, we designed the constraint prediction from point cloud (CstPnt) module to obtain MAD, Adj, PT.
-The CstPnt module is based solely on local features, experiments show that it exhibits strong generalization ability on unseen datasets, and its training data can be easily obtained by processing unlabeled B-Rep data.
-Following constraint acquisition, we separately extract features from MAD, Adj, PT, and apply attention mechanism to weight these features, thereby facilitating their effective utilization.
-We quantitatively evaluate CstNet on ABC [20] and Param20K dataset: constraints prediction, classification, and rotation robustness evaluation.
-For constraints prediction, our CstPnt module (in CstNet Stage 1) surpasses the state-of-the-art (SOTA) by around 5% and 14% instance accuracy in ABC (test set) and Param20K (unseen dataset), respectively.
-For classification and rotation robustness on Param20K, CstNet outperforms the SOTA by 3.52% and 26.17% instance accuracy, respectively.
+To mitigate the impact of point cloud unordered nature, we encode MAD-Adj-PT as point-wise features.
+As a result, each point with constraint representation in point cloud is defined as $(x,y,z,MAD,Adj,PT)$.
+The aforementioned processing transform the traditional constraint representation into deep learning-friendly form.
+After that, we designed the Constraint-Aware Feature Learning Network (CstNet), which consist of two stages.
+Stage 1 is designed to obtain MAD-Adj-PT.
+The architecture design is based solely on point cloud local features, experiments show that it exhibits strong generalization ability on unseen datasets, and its training data can be easily obtained by processing unlabeled BRep data.
+Following MAD-Adj-PT acquisition, Stage 2 separately extract features from MAD-Adj-PT, and apply attention layers to weight these features, thereby facilitating their effective utilization.
+We quantitatively evaluate CstNet on ABC [22] and Param20K dataset: MAD-Adj-PT prediction, classification, and rotation robustness evaluation.
+For MAD-Adj-PT prediction, CstNet surpasses the state-of-the-art (SOTA) by around 5% and 14% instance accuracy in ABC (test set) and Param20K (unseen dataset), respectively.
+For classification and rotation robustness experiments on Param20K, CstNet outperforms the SOTA by 3.49% and 26.17% instance accuracy, respectively.
 In summary, our contributions are threefold:
-  • We introduced a deep learning-friendly vectorized representation and designed the first constraint-aware feature learning network CstNet for effective constraint acquisition and leveraging for parametric point cloud.
-  • We built the first parametric-purpose Param20K dataset, which consists of about 20K shape instances of 75 classes.
-Each instance include mesh, point cloud, and B-Rep data.
-Given the current scarcity of labeled CAD shape datasets, Param20K offers a solid basis for future studies in parametric CAD domains.
+  • We introduced a deep learning-friendly constraint representation and designed the first constraint-aware feature learning network CstNet for effective constraint acquisition and leveraging for parametric point cloud.
+  • We built the first parametric-purpose Param20K dataset, which consists of about 20K CAD instances of 75 classes.
+Each instance include mesh, point cloud, and BRep data.
+Given the current scarcity of labeled CAD shape datasets [20, 16], Param20K offers a solid basis for future studies in parametric CAD domains.
   • We conducted extensive experiments on constraint acquisition, classification, robustness evaluation and ablation studies.
-The performance of CstNet has a significant improvement compared with SOTA methods which usually focus on geometric data only.
+The performance of CstNet has a significant improvement compared with SOTA which usually focus on geometric features only.
 2 Related Work
@@
 View-based methods typically project the point cloud from multiple views, the obtained figures are then fed in convolutional neural networks (CNNs) for feature extraction [44, 19, 9, 65, 64, 51].
-How to address information loss during the projection and optimizing viewpoints play a critical role.
+How to address information loss during the projection and optimizing viewpoints plays a critical role.
 Voxelization-based methods convert point clouds into 3D voxels, and then employ 3D CNNs for feature extraction [36, 57, 41, 49, 21, 24].
 The size and orientation of 3D voxels significantly influences the accuracy of these methods, and information loss occurs during the conversion to voxels.
-Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [36, 37, 27, 12, 57, 2, 43, 30, 64, 66, 11, 9, 55, 54, 53, 38, 31, 29, 46, 51, 5, 63].
+Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [38, 39, 29, 14, 62, 3, 47, 33, 69, 71, 13, 11, 60, 59, 58, 40, 34, 32, 50, 55, 6, 68, 7].
 The unordered nature and rotational invariance of point clouds make it challenging to employ methods similar to traditional CNNs for feature extraction, and kNN is a commonly adopted approach.
@@
 Other CAD-specific studies.
-B-Rep-based methods typically involve B-Rep data conversion and feature extraction, the challenge lies in converting B-Rep into deep learning compatible format.
-These approaches are commonly applied to classification [15], segmentation [21, 6], assembly [16, 48], engineering sketch generation [13], and CAD operation sequence prediction [50, 67].
-Mesh-based learning methods use triangle models as input, such as STL and OBJ, contributions often rely on mechanical datasets provided by the authors.
-These methods are primarily used for classification [32], retrieval [4], segmentation [44], and B-Rep generation [10].
-Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [10, 49, 34, 40, 28], 3D models [23, 24, 62], or CAD operation sequences [56].
+BRep-based methods typically involve BRep data conversion and feature extraction, the challenge lies in converting BRep into deep learning compatible format.
+These methods are commonly applied to classification [17], segmentation [23, 8], assembly [18, 52, 66], engineering sketch generation [15], and CAD operation sequence prediction [54, 72].
+Mesh-based methods use triangle models as input, such as STL and OBJ, contributions often rely on mechanical datasets provided by the authors.
+These methods are primarily used for classification [35], retrieval [5], segmentation [48], and BRep generation [12].
+Sketch-based learning takes free-hand sketches as input, with outputs typically being engineering sketches [12, 53, 37, 42, 31], 3D models [25, 26, 67], or CAD operation sequences [61, 45, 46].
 A key challenge is incorporating the regularity and connectivity of primitives, which is typically overlooked in graphic object analysis.
@@
 3.1 Is Constraint Important?
-We aim to utilize constraints to distinguish visually similar but functionally distinct CAD shapes.
 To answer whether constraints could enhance deep learning methods’ performance, we designed the following validation experiments.
-The objective is to compare the prisms shown in Fig.
+The objective is to discriminate prisms shown in Fig.
 2 with cuboids.
@@
 Consequently, eight independent experiments are conducted to cover all angles.
-As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them.
-The constraint-aware model is built on PointNet++ backbone, which predicts constraints and then adopts it for classification.
-The constraints are represented as the relation between the point attached primitive and a reference plane, more details in supplementary.
+As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it more challenging to distinguish them.
+The constraint-aware model is built on PointNet++ [39] backbone, which predicts constraints and then adopts it for classification.
+The constraints are represented as the relation between the point attached primitive and a reference plane, more details in $Suppl.$
 Figure 2: Prisms compared to cuboid.
@@
 Figure 3: Classification accuracy of validation experiments.
-However, the constraint learning method designed in the validation experiments is only applicable to simple scenarios.
+However, the constraint-aware learning method designed in the validation experiments is only applicable to simple scenarios.
 To develop a more generalized approach, we further investigated the constraint representation and deep learning architecture tailored for parametric point clouds.
 3.2 Constraint Representation
-In general, constraints should be defined as point-wise attribute for point clouds, analogous to normal vector.
+In general, constraint representation should be defined as point-wise attribute for point clouds, analogous to normal vector.
 As point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation.
@@
 However, once the constraint is applied, the positional relations between primitives become fixed.
-Therefore, if the type of primitives and their relations are determined, the constraints on CAD shapes can be defined comprehensively.
+Therefore, if the type of primitives and their relations are determined, the constraint representation on CAD shapes can be defined.
 The point-wise primitive type can be represented as the type of the primitive to which the point is attached, using one-hot encoding.
@@
 Considering that the direction of primitives’ main axis is unique, as shown in Fig.
-4, relations such as parallel and vertical could be derived from the direction of their main axes.
+4, relations such as parallel and vertical could be derived from their main axes.
 However, some dimensional relations can not, such as the cylinders’ radius equivalence.
-Thus, additional data is required to supplement this information.
 Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitations on the distance and size of primitives.
@@
 The edges are key indicators of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges.
-For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by two-channel one-hot vector.
-In summary, each composition of constraints in parametric point cloud is illustrated in Fig.
+For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by one-hot vector.
+In summary, constraint representation in parametric point cloud is illustrated in Fig.
 5, and each point with constraint is expressed as $(x,y,z,MAD,Adj,PT)$.
-Figure 4: Main Axis Direction of primitives. Plane: normal vector. Cylinder, cone, sphere: rotation axis. Main Axis Direction requires unique process, details are provided in Supplementary.
-We represent constraints by MAD, Adj, PT, is this approach sufficient to handle most constraints?
-For PT, commercial CAD systems typically classify primitives into plan, cylinder, cone, sphere, and free-form surface.
-Among these, plane, cylinder, and cone are primarily influenced by constraints, while spherical and free-form surfaces rely heavily on their geometric features.
-Therefore, considering plane, cylinder, and cone is sufficient for constraint representation.
-For sphere and free-form surface, the Stage 2 of CstNet also incorporate geometric features (feed in coordinates).
-MAD correspond to directional constraints (e.g.
-parallel and vertical), and MAD combined with PT represent dimensional constraints (e.g.
-distance and tangent).
+Figure 4: MAD of primitives. Plane: normal vector. Cylinder, cone, sphere: rotation axis. MAD computation requires unique process, details are provided in $Suppl.$
+Is MAD-Adj-PT sufficient to handle most constraints?
+For PT, commercial CAD systems typically classify primitives into plane, cylinder, cone, sphere, and free-form surface.
+Among these, plane, cylinder, and cone are the most common primitives, e.g., for a Trunk of ABC [22], plane: 69.63%, cylinder: 17.51%, cone: 6.99%, other : 5.87%.
+Besides that, constraints usually applied to regular primitives, while free-form surfaces rely heavily on geometric features.
+Therefore, considering plane, cylinder, and cone is sufficient for common constraint representation.
+For free-form surfaces, the CstNet Stage 2 also incorporates geometric features.
+MAD corresponds to directional constraints (e.g.
+parallel, vertical), and MAD combined with PT can represent dimensional constraints (e.g.
+distance, tangent).
 Through MAD and Adj, most common constraint types can be covered.
@@
 3.3 Constraint Acquisition
-The Stage 1 of CstNet is designed to obtain constraints, as shown in the left of Fig.
+The Stage 1 of CstNet is designed to obtain constraint representation, as shown in the left of Fig.
 7.
-When B-Rep data is available, the constraint computation from B-Rep (CstBRep) module is developed to obtain constraints.
-For the input B-Rep data, tessellation is performed by OCCT to generate a mesh model, followed by Poisson disk sampling [25] via pymeshlab to ensure a uniform point distribution and control the point count, if point cloud generation is not needed, this step could be skipped.
-For adjacency computation, we identify valid edges firstly.
+When BRep data is available, the constraint computation from BRep (CstBRep) module is selected.
+For the input BRep data, OCCT [2] performs tessellation to generate a mesh model, followed by Poisson disk sampling [27] via pymeshlab to ensure a uniform point distribution and control the point count.
+If point cloud generation is not needed, this step could be skipped.
+For Adj, OCCT identify valid edges firstly.
 An edge is considered invalid if the two adjacent primitives are smoothly connected along the edge.
-Searching valid edges is accomplished by OCCT.
-Next, we compute the distance from each point to all valid edges.
-A point is marked as near an edge if the minimum distance is below a threshold (set to $0.08\times ShapeArea/4\pi$ in this paper).
-For the primitive type and main axis direction of a point, we first determine which primitive it attached, OCCT is then used to analyze the primitive, providing both the primitive type and main axis direction.
-In the absence of B-Rep data, such as working with mesh or voxel models, it can be converted to point cloud, after which the CstPnt module can be used to predict constraints.
-Since each point’s constraints are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary.
-The constraint computation of a point $\boldsymbol{p}$ could be divided into the following steps: Step 1: Identify neighbor points around $\boldsymbol{p}$.
+Next, computing the distance from each point to all valid edges.
+A point is marked as near an edge if the minimum distance is below a threshold (we use $0.08\times ShapeArea/4\pi$).
+For the PT and MAD, OCCT first determine which primitive it attached, and then analyze the primitive, providing both the PT and MAD.
+In the absence of BRep data, such as working with mesh or voxel models, it can be converted to point cloud, after that the constraint prediction from point cloud (CstPnt) module can be used.
+Since each point’s MAD-Adj-PT are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary.
+The MAD-Adj-PT computation of a point $\boldsymbol{p}$ could be divided into the following steps: Step 1: Identify neighbor points around $\boldsymbol{p}$.
 Step 2: From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, store in $\boldsymbol{\mathrm{P_{valid}}}$.
@@
 Step 4: Based on $\boldsymbol{\mathrm{P_{valid}}}$ and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
-While the above process can be carried out using traditional methods, it is highly complex, which is why a deep learning model is employed to achieve the same.
+While the above process can be carried out by traditional methods, it is highly complex, therefore CstPnt is designed to achieve the same.
 The design of CstPnt module is inspired by the above steps, with details in Fig.
 7.
-For Step 1, the traditional approach is to use the k-Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere.
-However, for constraint computation, it is more effective to search for neighbors along the shape’s surface.
+For Step 1, the k-Nearest Neighbors (KNN) algorithm is commonly used, which identifies neighbors within a sphere.
+However, for MAD-Adj-PT computation, it is more effective to search for neighbors along the shape’s surface.
 This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig.
 6.
-To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary.
-For Step 2, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding, the relation between points is designed as embedding subtraction, as shown in Eq.
+To this end, we proposed a neighbor searching method SurfaceKNN, details in $Suppl.$ For Step 2, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding, as shown in Eq.
 1.
-Step 3 and Step 4 involve computing MAD and PT, both steps can be accomplished by MLPs.
-We employed a U-Net-like structure for better data utilization.
+Step 3 and Step 4 can be accomplished by MLPs, we employed a U-Net-like structure for better data utilization.
 $\boldsymbol{f}^{\prime}_{i}=\sum_{f_{j}\in\mathcal{N}_{i}}\rho\Bigl{(}\rm{MLP}\bigl{(}\mathrm{Q}(\boldsymbol{f}_{\mathit{i}})-\mathrm{K}(\boldsymbol{f}_{\mathit{j}})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{\mathit{i}}).$ (1)
-Where $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication.
+Where $\boldsymbol{f}_{\mathit{i}}$ is the feature of $\boldsymbol{p}_{i}$, $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication.
 $\mathrm{Q}(\boldsymbol{f})$, $\mathrm{K}(\boldsymbol{f})$, and $\mathrm{V}(\boldsymbol{f})$ are defined in similar structure by concatenating $\boldsymbol{f}$ with the positional encoding $\Delta$, and then feeding it into MLPs, where $\Delta$ is defined as $\Delta=\boldsymbol{p}_{i}-\boldsymbol{p}_{j}$.
 Figure 6: Comparison of KNN and SurfaceKNN. Blue point: center of neighbor search, red point: searched neighbors.
-The structure of CstPnt reveals that only local information is utilized, which aligns with the local nature of the constraint representation we designed.
-For no global information is incorporated and most CAD shapes consist of a limited set of primitives, such as plane, cylinder, and sphere, the CstPnt could generalize to unseen datasets if pre-trained on a large scale dataset.
-The training data is derived from unlabeled B-Rep data, as a result, our CstPnt enables the effective use of existing large scale unlabeled B-Rep datasets, making it possible to serve meaningful tasks.
+The structure of CstPnt reveals that only local information is used, which aligns with the local nature of the constraint representation we designed.
+Since no global information is incorporated and most CAD shapes consist of a limited set of primitives, such as plane, cylinder, and cone, the CstPnt could generalize to unseen datasets after pre-training on a large dataset.
+The training data is derived from unlabeled BRep data, as a result, our CstPnt enables the effective use of existing large unlabeled BRep datasets, making it possible to serve meaningful tasks.
 Figure 7: Overview of CstNet.
@@
 Stage 1 is designed for constraint acquisition.
-When B-Rep data is available, the CstBRep module is utilized to extract constraints; otherwise, use the CstPnt module.
+When BRep data is available, the CstBRep module is utilized to extract constraints; otherwise, use the CstPnt module.
 Stage 2 performs constraint feature learning, facilitating a deeper understanding of CAD shapes.
@@
 3.4 Constraint-Aware Feature Learning
-The Stage 2 of CstNet is designed to leverage constraints effectively, which extracts features from MAD, Adj, PT separately, and attention between features is then applied to update constraint feature, with details in Fig.
+The Stage 2 of CstNet is designed to leverage constraints effectively, which extracts features from MAD-Adj-PT separately, and attention between features is then applied to update constraint feature, with details in Fig.
 7.
@@
 All experiments are performed on GeForce RTX 4090 GPU.
-Param20K dataset Param20K contains 19,739 multi-modal instances categorized into 75 classes.
-Each instance includes B-Rep data, mesh, and point cloud.
-All B-Rep data are collected from the following three sources.
+Param20K dataset: Param20K contains 19,739 multi-modal instances categorized into 75 classes.
+Each instance includes BRep data, mesh, and point cloud.
+All BRep data are collected from the following three sources.
 More details in the supplementary.
@@
 The CAD shapes on TraceParts are collected from manufacturing companies.
-Sources 2 (24.12 %): Provided by a CAD company [^1] https://www.newdimchina.com/ that collaborates with our team.
+Sources 2 (24.12%): Provided by a CAD company [1] that collaborates with our team.
 These shapes are all mechanical components.
@@
 4.2 Constraint Acquisition
-The extraction of constraints from B-Rep data is accomplished by the CstBRep module, with results visualized in the Ground Truth column of Fig.
+Extracting MAD-Adj-PT from BRep data is accomplished by the CstBRep module, with results visualized in the Ground Truth column of Fig.
 8.
-By referring to the model row, CstBRep performs well in extracting MAD, Adj, and PT from B-Rep data.
-Predicting MAD, Adj, and PT from point cloud is achieved by the CstPnt module.
-The training set is sourced from 25 trunks of ABC [20], with each trunk containing 10,000 B-Rep files.
-We utilized CstBRep module to extract constraints as Ground Truth for model training.
+By referring to the model row, CstBRep performs satisfactorily in both MAD-Adj-PT extraction.
+Predicting MAD-Adj-PT from point cloud is achieved by the CstPnt module.
+The training set is sourced from 25 trunks of ABC [22], each trunk containing 10,000 BRep files.
+We utilized the CstBRep to extract constraints as Ground Truth for model training.
 After the training finished, the MSE loss of MAD is 0.0559, the prediction accuracy of Adj and PT reached 94.62% and 94.23%, respectively.
-In subsequent experiments, the above pre-trained CstPnt module is utilized to predict MAD, Adj, and PT with its weights frozen and no further updates.
-To ensure that only point cloud is used as input across all methods when comparing, even when B-Rep data is available, the CstBRep module will not be activated to obtain MAD, Adj, and PT.
+In subsequent experiments, the above pre-trained CstPnt is utilized to predict MAD-Adj-PT with its weights frozen and no further updates.
+To ensure that only point cloud is used as input across all methods when comparing, even when BRep data is available, the CstBRep module will not be activated to obtain MAD-Adj-PT.
 After CstPnt was pre-trained, we evaluated it on a ABC new trunk and Param20K dataset.
-The ABC new trunk was treated as the test set, while the Param20K was considered as an unseen CAD shapes dataset, more results on other public available CAD shapes datasets in supplementary.
-Since there is no existing model that directly predicts MAD, Adj, and PT, we modified the published CAD-specific methods [58, 41] by adding MLPs as prediction head to enable these predictions.
+The ABC new trunk was treated as the test set, while the Param20K was considered as an unseen CAD shapes dataset, more results on other public available datasets in $Suppl.$ Since there is no existing model that directly predicts MAD-Adj-PT, we modified the published CAD-specific methods [63, 43] by adding MLPs as prediction head to enable these predictions.
 The experiment results are shown in Tab.
-1, from which our CstPnt achieves the highest accuracy in predicting MAD, Adj, and PT.
+1, from which our CstPnt achieves the highest accuracy.
 Furthermore, from ABC to Param20K, all methods’ performance decreased, but our method dropped least accuracy.
@@
 Ours 0.1390 87.95 86.52
-Figure 8: Constraint prediction visualization on ABC and Param20K. MAD is processed for visualization, details in supplementary.
+Figure 8: Constraint prediction visualization on ABC and Param20K. MAD is processed for visualization, details in $Suppl.$
 4.3 Classification
-To validate the ability of our CstNet in identifying CAD shapes that are visual similar but different in constraints, we conducted experiments on the Prism-Cuboid datasets presented in the Problem Statement section Sec.
+To validate the ability of our CstNet in identifying CAD shapes that are visually similar but different in constraints, we conducted experiments on the Prism-Cuboid datasets presented in Sec.
 3.1.
 Results shown in Fig.
-9, in which our CstNet achieved best instance accuracy in distinguish 50°50°50  ° to 87°87°87  ° prisms and cuboids, which demonstrate the strong ability of CstNet in recognizing shapes with similar appearance but constraints different.
-For the binary classification of 89°89°89  ° prisms and cuboids, all methods achieved approximately 50% instance accuracy, this is attributed to the differences between them are so subtle that no one is capable to distinguishing them.
+9, in which our CstNet achieved the best instance accuracy in distinguishing 50°50°50  ° to 87°87°87  ° prisms and cuboids, which demonstrates the strong ability of CstNet in discriminating CAD shapes with similar appearance but constraints different.
+For the binary classification of 89°89°89  ° prisms and cuboids, all methods achieved approximately 50% instance accuracy, which is attributed to the differences between them being so subtle that no one is capable of distinguishing them.
 Figure 9: Classification results on Prism-Cuboid dataset.
 The classification of prisms and cuboids is predominantly influenced by constraints, while the comprehension of common CAD shapes necessitates the integration of both constraint and geometric features.
-CstNet achieves this by concatenating XYZ with constraint in Stage 2, thereby integrating geometric features into constraint features.
-To evaluate CstNet’s performance on common CAD shapes, we conducted experiments on Param20K dataset (more results on public CAD shapes dataset in Supplementary), results are presented in Tab.
+CstNet achieves this by concatenating xyz with MAD-Adj-PT in Stage 2, thereby integrating geometric features into constraint features.
+To evaluate CstNet’s performance on common CAD shapes, we conducted experiments on Param20K dataset, more results on other public available datasets in $Suppl.$ Results are presented in Tab.
 2, where our CstNet outperformed other methods across both metrics, validated the effectiveness of utilizing constraint and geometric features.
@@
 PointNet [38] 81.30 83.21 82.06 85.18
+PointConv [55] 82.30 85.62 84.91 87.92
 PointNet++ [39] 83.70 86.37 85.30 87.94
-PointConv [51] 82.30 85.62 84.91 87.92
+HPNet [63] 83.85 80.87 80.95 87.80
+ParSeNet [43] 83.96 79.50 78.52 85.66
+PTV2 [56] 85.14 73.80 72.49 89.19
 DGCNN [50] 85.40 87.28 86.43 89.17
 3DGCN [32] 86.42 88.25 87.63 89.81
+PTMamba [30] 86.45 87.28 86.63 91.34
 Ours 89.94 91.06 90.34 92.72
@@
 Since point clouds are pre-processed by aligning their centroids to the origin and scaling their ranges into a unit cube, the performance of all methods is unaffected by point cloud shifts and scaling.
-Comparison results depicted in Fig.
+Comparison results are depicted in Fig.
 10, from which our CstNet demonstrates strongest robustness to point cloud rotation.
@@
 4.4 Ablation Studies
-Our CstNet leverages all three components of constraint: MAD, Adj, and PT.
+Our CstNet leverages all three components of constraint representation: MAD-Adj-PT.
 Additionally, we utilized the SurfaceKNN to obtain more efficient neighbors for each point.
 To validate the effectiveness of our design, we conducted a series of ablation studies on Param20K dataset.
-For each component of constraint, we evaluated their importance by disabling one or all at a time.
-This operation is performed by setting the disabled components as a zero tensor before feeding into CstNet.
+For each component of constraint representation, we evaluated their importance by disabling one or all at a time.
+This operation is performed by setting the disabled components as a zero tensor.
 Classification results are presented in Tab.
-3, which show that all three components contribute to improving the performance of our CstNet.
+3, which show that all three components contribute to improving CstNet’s performance.
 To assess the effectiveness of SurfaceKNN, we compared it with KNN.
@@
 The results are presented in Tab.
-4, from which our method demonstrates best performance under both predicted and label constraints, indicating that our design achieves the best utilization of constraints.
-Furthermore, although the accuracy using predicted constraints is lower than that of using label constraints, it still shows a significant improvement compared to the baselines without constraints.
-This highlights the effectiveness of the predicted constraints in improving model performance.
-Table 4: Effects on Stage2 Backbone. Upper half: predicted constraints from CstPnt, lower: label constraints.
+4, from which CstNet demonstrates best performance under both predicted and label MAD-Adj-PT, indicating that our design achieves the best utilization of MAD-Adj-PT.
+Furthermore, although the accuracy of using predicted MAD-Adj-PT is lower than that of using label MAD-Adj-PT, it still shows a significant improvement compared to the baselines without MAD-Adj-PT.
+This highlights the effectiveness of the predicted MAD-Adj-PT in improving model performance.
+Table 4: Effectiveness of Stage2 Backbone. Upper half: predicted constraint representation from CstPnt, lower: label constraint representation.
 Backbone-S2 Acc Acc* F1 mAP
@@
 5 Conclusions
-In this paper, we studied the constraints of CAD shapes and introduced its deep learning-friendly expression.
-Afterward, we proposed the CstNet for extracting and leveraging constraints.
-Additionally, we built up a multi-modal CAD shape dataset Param20K.
-Extensive experiment results demonstrate that our CstNet achieves state-of-the-art performance on both accuracy and robustness.
-Comprehensive ablation studies further validate the effectiveness of constraints, SurfaceKNN, and network backbone.
+In this paper, we studied the constraints of CAD shapes and introduced its deep learning-friendly representation.
+Afterward, the CstNet is proposed for extracting and leveraging constraint representation.
+Finally, we built a multi-modal CAD dataset Param20K.
+Extensive experiments demonstrate that the CstNet achieves SOTA performance on both accuracy and robustness.
+Comprehensive ablation studies further validate the effectiveness of the proposed constraint representation, SurfaceKNN, and network backbone.
 CAD shapes possess numerous attributes.
-In this paper, we analyze the constraints from a functional perspective.
+We analyze the constraints from a functional perspective in this work.
 In future work, we plan to investigate additional aspects such as part design and manufacturing procedure, to achieve a deeper understanding of CAD shapes.
-Supplementary Material
-F Overview
-This supplementary material provides additional details that supplement the main paper and includes more experimental results.
-In Sec.
-G, we present the details of network architecture and data preparation in validation experiments.
-In Sec.
-H, we provide the data construction and static of Param20K dataset.
-Finally, Sec.
-I offers data process for visualization and more results of constraint acquisition and classification experiments.
-In Sec.
-J, we show more details of our constraint representation and constraint learning methods.
-G Validation Experiment
-G.1 Constraint Expression
-This section introduces the constraint expression in the validation experiments.
-The constraint expression in validation experiments is defined in a simple manner: for a point, its constraint is defined by the relationship between the plane it attached and the reference plane, as shown in Fig.
-11.
-This calculation process is carried out by openCASCADE Technology (OCCT).
-The constraint is represented in a one-hot encoding, where $(\mathit{1,0,0})$ indicates perpendicular, $(\mathit{0,1,0})$ indicates parallel, and $(\mathit{0,0,1})$ indicates other constraints.
-Each point in the point cloud is expressed as $(Coordinate,constraint)$.
-For example, if the coordinate of a point is $(0.1,0.2,0.3)$, and the plane it attached is perpendicular to the reference plane, it is expressed as $(0.1,0.2,0.3,\mathit{1,0,0})$, where $(0.1,0.2,0.3)$ represents the point’s coordinate, and $(\mathit{1,0,0})$ denotes the type of constraint.
-Figure 11: Constraint expression in validation experiments. The red point’s constraint is expressed as the constraint type between the magenta plane and the reference plane, using one-hot encoding.
-G.2 Dataset Generation
-This section introduces the dataset generation procedures in the validation experiments.
-The dataset generation for the validation experiments consists of the following steps:
-  1. 1.
-    Building the parametric templates for cuboids and prisms (with angles of 50°50°50  °, 60°60°60  °, 70°70°70  °, 80°80°80  °, 82°82°82  °, 85°85°85  °, 87°87°87  °, 89°89°89  °).
-Afterwards determining the dimension range for each parameter.
-The parametric templates and the dimensions are shown in Fig.
-12.
-  2. 2.
-    Assigning random values to the parameters in parametric templates. Each parametric template is instantiated into 5,000 B-Rep format CAD shapes, which are stored in STEP files [35].
-  3. 3.
-    Using OCCT to read the STEP files and convert them into meshes. Poisson disk sampling is then applied to sample points from meshes.
-  4. 4.
-    For the sampled points, OCCT is used to calculate the plane to which each point is attached.
-The constraint type for each point is then determined based on the relationship between the attached plane and the reference plane.
-  5. 5.
-    Rotating the generated point clouds randomly along the XYZ axes, within a range of [25°-25°- 25  °, 25°25°25  °].
-  6. 6.
-    Assigning the generated point clouds to each experiments.
-The validation experiments involve binary classification of cuboids and prisms with specific angles.
-The point clouds used in each of the eight experiments are as follows (80 % for training, and 20 % for testing):
-    • Experiment 1: Cuboid $\times$ 5000, Prism 50°50°50  ° $\times$ 5000
-    • Experiment 2: Cuboid $\times$ 5000, Prism 60°60°60  ° $\times$ 5000
-    • Experiment 3: Cuboid $\times$ 5000, Prism 70°70°70  ° $\times$ 5000
-    • Experiment 4: Cuboid $\times$ 5000, Prism 80°80°80  ° $\times$ 5000
-    • Experiment 5: Cuboid $\times$ 5000, Prism 82°82°82  ° $\times$ 5000
-    • Experiment 6: Cuboid $\times$ 5000, Prism 85°85°85  ° $\times$ 5000
-    • Experiment 7: Cuboid $\times$ 5000, Prism 87°87°87  ° $\times$ 5000
-    • Experiment 8: Cuboid $\times$ 5000, Prism 89°89°89  ° $\times$ 5000
-Figure 12: Parametric templates for prisms and cuboids. $a,b,c\in[0.3,2.3]$, for prisms: $b>\frac{c}{tan\alpha}$.
-G.3 Model Design
-This section introduces the structure of the constraint-aware model used in the validation experiments.
-The constraint-aware model first employs a PointNet++ [37] backbone to predict point-wise constraints.
-These constraints are then concatenated with the point coordinates and passed through another PointNet++ backbone to determine the point cloud category.
-The model’s loss function is the sum of the classification loss and the constraint loss, as shown in Fig.
-13
-Figure 13: Constraint-aware model structure.
-H Param20K Dataset
-H.1 Why build up the Param20K dataset?
-The limited availability of B-Rep datasets poses barriers to deep learning on CAD shapes.
-Most existing CAD shape datasets consist of mesh files, such as MCB [18] and ESB [14].
-While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information.
-In contrast, B-Rep data [21, 61] serves as the native representation of CAD shapes and is therefore more suitable for dataset construction.
-However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files.
-Although the ABC [20] contains a large number of B-Rep files, it remains unlabeled.
-H.2 Dataset Construction
-This section provide details of purification process and data acquisition of Param20K dataset.
-Purification: The purification process involved removing broken, duplicated, or overly complex CAD shapes requiring more than 4,000 points for point cloud representation.
-Point cloud: The point cloud data in Param20K dataset containing about 2,200 points, point constraints is also included, each point is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency), the meshes and point clouds are generated from B-Rep data.
-H.3 Dataset Statics
-This section presents the shapes count and proportion statistics for each category in the Param20K dataset. Additionally, shapes from each category are showcased.
-Param20K dataset is divided into 80 % for training and 20 % for testing.
-The class distribution is shown in Fig. 14.
-The proportion of shapes in each category are shown in Fig. 15.
-Shapes instantiated from the parametric templates designed by our team names begin with ”Num”.
-Example shapes from each category are shown in Fig. 16.
-Figure 14: Quantity distribution of Param20K dataset.
-Figure 15: Shapes proportion of Param20K dataset.
-Figure 16: Example shapes for each category in Param20K dataset.
-I Constraint Acquisition and Classification Experiments
-I.1 Process to Visualize Main Axis Direction
-The main axis direction of a point, denoted as $v=(x,y,z)$, is visualized as $R=\frac{x+1}{2}$, $G=\frac{y+1}{2}$, and $B=\frac{z+1}{2}$.
-Since the main axis direction is represented as a unit vector, $-1\leqslant x,y,z\leqslant 1$, which results in $0\leqslant\frac{x+1}{2},\frac{y+1}{2},\frac{z+1}{2}\leqslant 1$.
-I.2 Constraint Visualization of CstBRep
-More constraints extracted by the CstBRep module from the B-Rep are visualized in Fig. 17.
-Figure 17: More constraint extraction results from the B-Rep data. ADJ: Adjacency, PMT: Primitive Type, MAD: Main Axis Direction.
-I.3 Constraint acquisition on MCB
-We evaluated the pre-trained CstPnt module and ParSeNet [41], HPNet [58] on the CAD shape dataset MCB [18] to predict constraints.
-For the MCB consists of mesh files, it was not possible to extract Ground Truth constraints for comparison, therefore no ground truth presented in Fig.
-18, and no numerical results are available.
-Based on the results visualized in Fig.
-18, CstPnt demonstrated a high level of constraint prediction accuracy even on an unseen dataset.
-Figure 18: Constraint prediction on MCB.
-I.4 Classification on MCB
-We conducted evaluations on the CAD shape dataset MCB [18], with results presented in Tab.
-5, from which our CstNet achieved the highest scores across all metrics, conclusions similar with experiments on Param20K dataset.
-Table 5: Classification results on MCB. Acc: accuracy over instance %, Acc*: accuracy over class %, F1: F1-score, mAP: mean average precision %.
-Method Acc Acc* F1 mAP
-PointCNN [27] 93.89 81.85 83.86 90.13
-PointNet [36] 86.78 67.70 86.55 74.08
-PointNet++ [37] 87.45 73.68 88.32 91.33
-SpiderCNN [57] 93.59 79.70 81.30 86.64
-PointConv [51] 93.25 80.24 71.31 82.19
-DGCNN [46] 92.54 74.47 76.12 74.27
-3DGCN [29] 93.71 78.71 84.59 84.35
-Ours 96.87 89.21 89.85 93.17
-J Constraint Learning Method Details
-J.1 Defects of General Constraint Representation
-The simple constraint expression in validation experiments is insufficient for a comprehensive definition, as selecting an appropriate reference plane for complex CAD shapes is challenging.
-Moreover, traditional constraint representations suffer from non-uniqueness, as shown in Fig.
-19, defining the relations between three planes involves multiple constraint options, therefore, this approach is not well-suited for deep learning models.
-Figure 19: Non-uniqueness of traditional constraint expression. Left: Double vertical constraints, right: Vertical and parallel constraints.
-The constraint is defined as a graph representation in some circumstances, where each node represents a primitive, and the edge between nodes represent their constraints.
-However, this approach is not well-suited for point cloud analysis.
-First, in cases where only point clouds are provided as input, the number of primitives is hard to obtain.
-Second, due to the unordered nature of point cloud, it is difficult to associate points with the corresponding primitives, making label design very challenging.
-Additionally, the graph representation essentially records the constraint relationships between each pair of primitives, for the CAD shapes contain many primitives, recording the relative positions between every pair of primitives would result in an overwhelming amount of data, making the expression of these relative positions impractical.
-J.2 Unique Process of MAD
-This section provides the process of Main Axis Direction and example constraint representation.
-The main axis direction is processed to ensure its uniqueness, as one axis can determine two unit vectors with opposite direction.
-For a given axis, we first randomly select one of the two vectors, denoted as $v(x,y,z)$, and then process it using the following algorithm:
-Data: Unit vector $v(x,y,z)$
-Result: Unit vector after direction unification
-1 if $z<0$ then
-2       $v=-1\times v$;
-3       Return $v$;
-4
-5else if $z==0$ then
-6       if $y<0$ then
-7             $v=-1\times v$;
-8             Return $v$;
-9
-10      else if $y==0$ then
-11             if $x<0$ then
-12                   $v=-1\times v$;
-13                   Return $v$;
-14
-15
-16
-return $v$
-A point with constraint is expressed as Fig. 20.
-Figure 20: Expression of a point with constraint.
-J.3 CstBRep Module
-This section explains why it is necessary to compute valid edges in the CstBRep module.
-The CstBRep module is used to compute constraints from B-Rep data, and the computation process consists of the following steps:
-Step 1: Point cloud generation (optional).
-Step 2: Calculation of valid edges and the distance from each point to these valid edges. If the minimum distance is below a threshold, the point is classified as a edge-near point.
-Step 3: Calculation of the primitives to which each point attached, followed by analysis to determine point’s Primitive Type and the Main Axis Direction.
-In Step 2, valid edges are computed, as this helps to reduce the potential for data to mislead the deep learning model.
-The definition of valid edge is illustrated in Fig.
-21.
-This definition is based on the observation that a complete cylinder in the B-Rep data stored in STEP files is often split into two half-cylinder faces.
-As a result, two generatrices form edges on the cylindrical surface, but the deep learning model cannot predict these edges, as shown in Fig.
-22.
-Furthermore, in other cases, edges corresponding to smoothly connected faces are also difficult to identify, leading to confusion for deep learning models.
-Figure 21: Zebra stripes of primitives corresponding to invalid and valid edges. Left: G1 or higher continuous at the invalid edges, right: G0 continuous at the valid edges.
-Figure 22: Cylinder in B-Rep representation.
-A complete cylindrical surface is represented as a combination of two half-cylinder surface, resulting in edges along the generatrices that cannot be recognized, thereby causing confusion for deep learning model.
-J.4 CstPnt Module
-This section provides a detailed explanation of the valid points in the computational steps that inspire the design of the CstPnt Module.
-The CstPnt Module is used to predict constraints from point clouds. The design of the CstPnt Module is inspired by the following calculation process:
-For a point $\boldsymbol{p}$ in point cloud, its constraint could be calculated by the following steps:
-  1. 1.
-    Identify neighbor points around $\boldsymbol{p}$.
-  2. 2.
-    From all neighbor points, identify those that belong to the same primitive as $\boldsymbol{p}$; these points are referred to as valid points.
-  3. 3.
-    Fit shapes such as cylinders or planes using valid points, and determine the primitive type with the smallest fitting error.
-  4. 4.
-    Based on the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
-In these steps, it is necessary to compute the valid points.
-We define the valid point as the neighbors that lie on the same primitive as the center point $\boldsymbol{p}$, as visualized in Fig.
-23.
-These valid points can be used to determine whether the center point is near an edge, as well as to calculate the main axis direction and primitive type of the attached primitive, while the invalid points do not contribute to these calculations.
-Figure 23: Valid and invalid points for constraint calculation.
-J.5 SurfaceKNN
-This section introduces the algorithm of SurfaceKNN.
-SurfaceKNN based on the assumption that when KNN is applied with a small number of neighbors, the output points are searched along the shape’s surface.
-This assumption has been validated in most cases.
-Leveraging this assumption, SurfaceKNN is accomplished by applying KNN with small number of neighbors iteratively.
-During each iteration, KNN is used to search a small set of neighboring points, which then serve as new central points for subsequent searches.
-This process continues until the desired number of points is obtained, as illustrated in Fig.
-24.
-Figure 24: Algorithm of SurfaceKNN.
+Acknowledgements
+This work was supported by National Key Research and Development Program of China (Grant No.2022YFB3303101).